A Bibliography of Publications in Science and Public Policy: 2020–2029

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

15 May 2024 Version 1.12

Title word cross-reference

19 [Boz22, FLBM23, KLL23, NNS^{+}22, RML^{+}21, SLSC^{+}21, SWAD24, WvSG22, XLL22, YAB^{+}20]. 1970s [RAB^{+}23].

2020 [EG20]. 21st [MB20].

absorptive [CBUS23, dSTRC20, VJMHdLNA20]. academia [HL22, JGK22]. Academic [GCGAMASG21, KKK^{+}21, SU22, ZS21, dJB22, Abd23, Bre22, Cha22, Dui20, GRFBL21, HR20, LMS21, MSF^{+}22, OI22, SFvD^{+}21, Woe21, Zou22, dJBN22]. academics [AOF22, Nel24, SU22]. academies [Bav23]. Accelerating [YKW21]. acceptance [SC21b]. accountability [RKT23]. Achieving [Bre22]. acquisition [CBUS23]. across [CNA^{+}20b, CNA^{+}20a, CNA^{+}20c, JWT^{+}23]. Actionable [DC23]. actors [AYDQ^{+}21, JS23]. adaptation [HL20]. Addressing [SGW21, Mar23]. administrative [Woe21]. Adopting [TTF20]. adoption [ABC^{+}21, NRSA23, OSV^{+}23]. advantage [Her23]. Advice [AAW^{+}20, Bla21b, Bla21a, DC23]. advisory [GWS23, SWAD24]. advocate [SHK^{+}20]. aeronautical [CBUS23]. affairs [Nel24]. affect [Pan20, YH21]. affiliations [HL22]. affordances [SU23]. Africa [AGKL21, Kru20, TBM21]. against [CK21]. age [DS23]. agencie [Bar23]. agencies [Abd23, HV23, SMPB24]. Agency [Loy22]. agenda [Par20, SPB20]. agenda-setting [SPB20]. Agendas [AAW^{+}20]. agreements [ASFB^{+}23, WS23]. AI [CP23, CLvH22]. AIDS [XHQ^{+}22]. Air [CBUS23]. alienation [Lil20]. alignment [SCY21]. allocated [Phi22]. allocation [GRL23, HTB22, ZXW^{+}20]. alongside [hPKvR22]. Alternative [WWF^{+}20]. Ambiguity [PEG20]. ambitions [Mar23]. ambivalence [Ano22, wY22]. America [EKFR^{+}23, Ter23]. among [CH20a, CH20b, KKK^{+}21, Nel24, SFvD^{+}21, ZYH^{+}20]. Analysis [AMDN21, KLL23, MY21, ABM21, BMCF21, Cha22, HYO^{+}21a, HYO^{+}21b, IW20, LLM^{+}22, MVNSdUC22, REL22, SC22, XG23, ZG22, dlGGM21]. Anti [NN22]. Anti-transparency [NN22]. antibiotic [BWG22, BGW22]. antibodies [dSFSN23]. anticipation [SFvD^{+}21]. any [LMS21]. apart [Clo21]. Application [Bor21, RK23, TTF20]. Application-oriented [Bor21]. applications [MA22]. Applied [CM21, Cav24, CHH^{+}24]. appraisal [EOS24]. approach [AYDQ^{+}21, DWK23, EKFR^{+}23, KLL23, LLRD23, PBW20]. approaches [SBH^{+}22]. Approaching [Par20]. appropriability [ORZJU24]. appropriation [Bar23]. arbiter [SHK^{+}20]. Area [SC22]. art [OI22]. Artificial [ZXW^{+}20, AY23, OSV^{+}23, PM20]. Assessing [OF22, GKR^{+}21, Lau20, WB22]. Assessment [MVNSdUC22, ABS21, Bla21b, Bla21a, HV23, LSvH22]. assessments [ASGRM23, DSW23]. asset [Tas24]. assets [ORZJU24]. assurance [FKF^{+}23]. asymptomatic [NNS^{+}22]. Atlantic [Pol23]. attention [MB20]. attentional [KKK^{+}21]. attitudes [OSV^{+}23, Pan20]. attributes [MG22]. Australian [Woe21]. Austria [RO24]. autonomy [BL22].

Balance [Lag21]. Balanced [Bor21].

Balancing [FKF^{+}23, Kru20]. barrier [Mar23]. Barriers [dMSVF23, Don20, LLRD23]. based [CZtB21, FZW22, NNS^{+}22, PEG20, Tas24]. Bavaria [FLBM23]. Bay [SC22]. be [CP23]. bedside [SHRA20]. behavior [BS22a]. behind [HL22, RP23]. belonging [HSAR23]. bench [SHRA20]. beneficiary [CBUS23]. Benefits [DP21]. Bermuda [Nik24]. best [BTT22]. Between [WLW20, Clo21, EKFR^{+}23, FMF20, FKF^{+}23, HYO^{+}21a, HYO^{+}21b, JGK22, LLRD23, PDM23, SHRA20, SCY21, ZAB^{+}20, ZAB^{+}21, dRS20]. beyond [TCWZ21]. bias [PV24]. bibliographic [MVNSdUC22]. bibliometric [MSF^{+}22, ZG22, Zou22]. bibliometrics [BSTG21]. Big [Vin22]. biodiversity [VPK20]. bioeconomy [ALK22, SBH^{+}22, Sol21]. biomedical [SHRA20]. Biometrics [Law24]. Biotech [YXX21]. biotechnology [HÅ23]. blockchain [JLX21, LCL23]. Blood [NLOL24]. boards [GWS23]. bodies [SWAD24]. Boosting [Ehs21]. bottlenecks [Sil24]. Boundary [AVZ21, Hen23, HSSS21, Ano23a, Ano23d, EET22, JWT^{+}23, MCDD23, OI22, SHK^{+}20, SPB20, TG22]. boundary-object [EET22]. Brain [MY21]. Brain-Drain [MY21]. Brazil [Don20, TRdRABS22, dSFSN23]. Brazilian [AVZ21, CBUS23, MG22]. Bred [MY21]. Bridging [Ano23a, CCF24, EET22]. Bringing [NLOL24]. Broadening [Bav23]. broader [ASGRM23, OF22, WB22]. broker [SHK^{+}20]. brokering [Ano23d, MCDD23]. brokers [Bav23]. Building [ZAB^{+}20, ZAB^{+}21, hPKvR22, RO24]. bureaucratization [Woe21]. Business [LSvH22, LLM^{+}22].

calculation [JS23]. Calibrating [GST^{+}20]. Can [BS22a, XG23, HTB22]. Canada [AKC22, DF22]. Canadian [ZAB^{+}20, ZAB^{+}21]. capabilities [hPKvR22, SMPB24]. capacity [CBUS23, dSTRC20, VJMHdLNA20]. capital [GCGAMASG21]. Capturing [Rud22]. cardiometabolic [vdKYYWR23]. Care [SGW21]. career [CHH^{+}24]. carry [BWMH22]. Case [AAW^{+}20, AVZ21, DP21, AIIOL22, CPT^{+}24, HLM^{+}23, HJMX23, Kru20, Loy22, Mo23, NLP23, NJV23, PDM23, PV24, QS23, RRC^{+}23, SWAD24, SX19, SX20, Sol21, SU22, ZS21]. cases [CK21, Don20]. catching [NCTd20]. catching-up [NCTd20]. category [PFPV20]. cause [Haw20]. causes [SFP23, ZXW^{+}20]. censorship [FKF^{+}23]. Centers [Lin21, XHQ^{+}22]. central [Pen23]. centre [HÅ23]. Centres [Ter23]. centric [CCT21a, CCT21b]. century [MB20]. CERN [Ano22, wY22]. certification [CPT^{+}24, KCL20, LXLY20]. Certifying [CPT^{+}24]. CGE [HYO^{+}21a, HYO^{+}21b]. chain [LCL23, WJZ20]. chains [MGM^{+}21, Qia23]. Chairs [HR20]. challenge [HL20]. Challenges [DP21, RML^{+}21, VPK20, YXX21, AJG22, BS22b, HJMX23]. Change [Ano23a, CLiM^{+}24, Don20, EET22, JWT^{+}23, Loy22]. changes [AYDQ^{+}21, CCT21a, CCT21b]. Changing [HLS20, AKC22]. Channels [FMF20, LKZC22]. characteristics [FS20, PBSA20]. chemical [MB20]. China [AJG22, CBWJ20, CZL^{+}20, CZtB21, HLM^{+}23, JLX21, JWFY22, LXLY20, LL22, Pan20, SX19, SX20, SGDW23, TCWZ21, WS23, XG23, YCL22, ZXW^{+}20, ZG22, ZFCL21, ZYH^{+}20]. China-style [ZFCL21]. Chinese [FZW22, MY21, WJZ20, WHH24, YLGZ23, ZQH^{+}23]. choice [Nel24]. circle [VJMHdLNA20]. citations [NCTd20]. cities [FZW22]. city [SX19, SX20]. civil [BWMH22]. Climate [Ano23a, BMA23, EET22, JS23, Loy22, Mar23, RP23]. Clinical [MRB20, RAM22]. clustering [KM23]. clusters [HJMX23]. Co [AGKL21, GFF^{+}23, LBCG^{+}23, MG22, NNAH22, RMP23, VBK^{+}23, Ano23b]. Co-creating [LBCG^{+}23, Ano23b]. co-creation [RMP23]. Co-design [GFF^{+}23]. co-inventions [NNAH22]. co-patents [MG22]. Co-production [AGKL21, VBK^{+}23]. Coevolution [PDM23]. collaboration [AJG22, BSTG21, CZL^{+}20, JGK22, LC23, MBT20, NNAH22, XHQ^{+}22]. collaborations [EKFR^{+}23, OI22, PBW20]. Collaborative [DP21, VV22, Bar23, Dui20, EG20, FZW22, GST^{+}20, PSK^{+}23]. collateral [TCWZ21]. collectives [Coh22]. Coloniality [Pol23]. CoM [Ano23a, EET22]. combat [BWG22, BGW22]. combating [AY23]. commitment [BL22]. commons [Sid20]. communication [FKF^{+}23]. communities [AGKL21, ABC^{+}21, HSAR23, JWT^{+}23]. community [FvL22, hPKvR22]. companies [CBUS23, KCL20]. comparative [CK21, SC22, ZQH^{+}23]. comparing [Bla21b, Bla21a]. comparison [Bre22, HHL21, TO21]. competence [SCY21]. Competing [HHL21]. Competition [HJMX23, ZFCL21]. complement [SGDW23]. complementary [ORZJU24]. Complexity [WDJ^{+}21, CPT^{+}24, Qia23]. compliance [MRB20]. comprehensive [ZQH^{+}23]. Computational [HSSS21, LLRD23]. conceptions [JGK22]. concepts [GRS24]. conceptual [AY23]. conceptualization [EOS24]. concern [Haw20]. conditional [XG23]. Conditions [FKS^{+}21, LKZC22, SU23, dJBN22]. conflicting [ALK22, CCF24]. conflicts [Bav23]. congress [Bla21a, Bla21b]. consequences [GFF^{+}23]. considering [HYO^{+}21a, HYO^{+}21b, LCL23]. constrain [Dui20]. Constructing [RML^{+}21]. construction [Coh22, PM20]. consumer [ST21]. Context [SPB23, GV20, HSAR23, SGW21]. context-development [GV20]. contexts [MP23]. contingent [CBUS23]. contrast [AD20]. conventional [KH22]. convergence [KCL20]. cooperation [ASFB^{+}23, MVNSdUC22, SDM23, vWBT20]. Cooperative [GSR22]. cooperatives [ZYH^{+}20]. coordination [HLM^{+}23, JWT^{+}23]. Coping [SLSC^{+}21]. coproduction [APGL20]. coronavirus [GZG^{+}23]. Corporate [MPCAC20, JWFY22, WZ23]. Corps [YKW21]. Correction [Ano22, Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, CNA^{+}20b]. Corrigendum [HYO^{+}21b]. corruption [AY23]. cost [GRFBL21, REL22]. could [Bav23]. Council [SWAD24]. councils [CM20]. count [JS23]. countries [RK23]. country [ABM21, SFQdM24, dSTRC20]. counts [Wil22]. coupling [MVNSdUC22]. COVID [Boz22, FLBM23, KLL23, NNS^{+}22, RML^{+}21, SLSC^{+}21, WvSG22, XLL22, YAB^{+}20, SWAD24]. COVID-19 [Boz22, FLBM23, KLL23, NNS^{+}22, RML^{+}21, SLSC^{+}21, WvSG22, XLL22, YAB^{+}20, SWAD24]. CREATE [BTT22]. Creating [AAW^{+}20, Ano23b, LBCG^{+}23]. Creation [AVZ21, RMP23]. credibility [BMCF21]. crises [GSR22, Sil24]. crisis [BWG22, BGW22]. CRISPR [SJ24]. criteria [FS20, Woe21]. Criterion [WB22]. critical [BBAK21, EOS24, Rud22, SCY21]. criticisms [FvL22]. Crops [NK21]. Cross [ABM21, TO21, ZX20]. Cross-country [ABM21]. cross-national [ZX20]. cross-regional [TO21]. cultural [PFPV20]. culture [DC23, NNS^{+}22]. cultures [Clo21]. cumulative [Her23]. curiosity [Ano22, wY22]. Cyborg [ISS22]. cycles [FZW22].

damage [TCWZ21]. Danish [HLS20]. data [DS23, MG22, MRB20, MPCAC20, REL22, Sid20]. data-driven [DS23, Sid20]. decarbonisation [Abd23]. decision [LKZC22]. decision-making [LKZC22]. decisions [YH21]. decomposition [MPCAC20]. Deconstructing [MA22]. Decoupling [TCWZ21]. definition [CM20, GCGAMASG21]. deliberation [WBM22]. Delphi [RR20]. demand [NRSA23]. Democracy [WDJ^{+}21]. Democratic [WvSG22]. Departments [Lag21]. dependence [Pan20]. design [BS22b, GFF^{+}23, KKS20, Mou24, SMPB24]. Designing [LLM^{+}22]. deteriorating [TCWZ21]. Determinants [SJ22, FS20, NPZ21, dSTRC20]. developed [dRS20]. developing [Fli22, GFF^{+}23, SFQdM24, dSTRC20, VPK20, dRS20]. Development [Lin21, CZtB21, GV20, Pen23, Pes21, TTF20, vWBT20]. developments [BCFU23]. devices [Woe21]. dialogue [SC21b]. differ [Ter23]. difference [LMS21, XG23]. difference-in-differences [XG23]. differences [XG23, ZYH^{+}20]. Different [CZL^{+}20, MPCAC20, NCTd20]. differentiation [Cav24]. differently [Par20]. diffusion [Hen23, HL21]. digital [CCF24, Hen23, PH23, SGW21, TSS23, VSGMF21]. dilemmas [Lau20]. Dimensions [PFPV20]. diplomacy [EKFR^{+}23, Fli22, KR23, Pol23, RAB^{+}23, RK23, RRC^{+}23, WS23]. discipline [HLL^{+}23]. discipline-specific [HLL^{+}23]. Disciplining [hKZ22]. disclosures [LS20]. Discontinuation [YKW21]. discourse [RMP23]. discovering [GRS24]. discrepancy [KKK^{+}21]. Discursive [Yeh21]. Discussion [QS23]. disease [GZG^{+}23, YLMYY22]. diseases [YLMYY22]. Disentangling [HSAR23]. disillusionments [ALK22]. distribution [CLvH22, GRFBL21]. Distrust [LRS24]. diverse [HTB22]. Diversification [BCFU23, AIIOL22, TO21]. diversity [GRL23, HSYW21, KH22, SGW21]. Do [APV21, DSW23, YH21, CCT21a, CCT21b, SGDW23, Ter23, Vin22]. Does [ASFB^{+}23, APV21, LXLY20, LMS21, WZ23, Yan21, Don20, FMF20, Kis20, Pan20, PFPV20, ZX20]. domestic [GSR22]. Dominican [GV20]. Drain [MY21]. Drawing [NK21]. drive [WBM22]. driven [DS23, FD24, Sid20, THPT20]. driver [TO21]. Drivers [ORZJU24, AOF22]. driving [SC21b]. drug [XLL22]. during [NNS^{+}22, SLSC^{+}21, WvSG22, ZQH^{+}23]. dynamic [SMPB24]. Dynamics [ALK22, Bar23, HLM^{+}23, Nel24, YLMYY22, ZW20].

E-hailing [SX19, SX20]. e-waste [AYDQ^{+}21]. each [ZG22]. early [RAB^{+}23]. eco [ORZJU24]. eco-innovation [ORZJU24]. ecological [RP23]. ecologies [ZS21]. Economic [LL22, GV20, HYO^{+}21a, HYO^{+}21b, SWA^{+}21]. economies [AY23]. economists [YAB^{+}20]. economy [Ben22, Sid20]. ecosystem [SFQdM24, VPK20]. ecosystems [NLP23]. Edited [NK21]. editing [SJ24, WS21]. Education [LER^{+}22, ABM21, Her23, TCWZ21]. Effect [WJZ20, LL22, MBT20, ZXW^{+}20]. Effective [Bla21b, Bla21a, RML^{+}21]. effectiveness [AD20, YLGZ23, ZX20]. Effects [ON21, SFvD^{+}21, TBM21, AOF22, BZG21, CZL^{+}20, SK23, THPT20, VJMHdLNA20, YCL22]. efficiency [ABM21, CZL^{+}20, FZW22, KCL20, TS23, WJZ20, ZFCL21, ZYH^{+}20]. efficient [HTB22]. efforts [Boz22, JS23]. embedded [Dui20]. Embracing [VBFLP22, Ano23c]. emerge [ZG22]. emergence [CBWJ20]. emergency [XLL22]. emerging [AY23, RAJV21]. Emissions [Mo23]. Empirical [IW20, ON21, BFH^{+}20, FZW22, ZX20]. Empowering [BS21]. enable [Dui20]. enabling [HL20, Rea22, dJBN22]. enactment [BMA23, Pra20]. end [Ano23b, LBCG^{+}23]. end-user [Ano23b, LBCG^{+}23]. Enemies [RAJV21]. Energy [Yeh21, JS23, KH22, Rud22]. Engage [Lag21]. engagement [AOF22, Coh22, CLiM^{+}24, EKFR^{+}23, GFF^{+}23, LMR^{+}20, SU22, ZS21]. Engaging [YXX21]. enhancement [ISS22]. enlargement [XHQ^{+}22]. entanglement [FE20]. enterprise [LXLY20]. enterprises [Ehs21, GSR22, SJ22]. entrepreneur [DSW23]. entrepreneurial [Coh22, MP23, QS23, SFQdM24]. entrepreneurs [Ano23d, MCDD23, MP23]. entrepreneurship [CK21, CLiM^{+}24, Hem23, KKK^{+}21, Rea22]. environment [EZI20, Lin20, THPT20]. Environmental [BZG21, Mo23, Don20, LL22, YH21, Loy22]. epistemic [FE20, JGK22, KM23]. equality [CPT^{+}24, SPB23, VV22]. era [VSGMF21]. Erratum [Bla21b, BWG22, CCT21a, CH20a, SX20]. ethical [BBAK21, ISS22, WS21]. ethics [ISS22, Nik24]. EU [Cav24, NN22, Nik24, VPK20]. Europe [Ano23e, CNA^{+}20b, CNA^{+}20a, CNA^{+}20c, EKFR^{+}23, HJMX23, MA23, MP23]. European [AMD24, AFF20, CPT^{+}24, Clo21, HLL^{+}23, LM23, PV24, PBSA20, RMP23, SC21a, TSS23]. evaluate [SHRA20]. Evaluating [CM21, JLX21]. Evaluation [ASGRM23, FS20, MA22, RCK23, SPB23]. events [HR20]. Evidence [HL21, JWFY22, KKK^{+}21, Mar23, NNAH22, WJZ20, XLL22, YCL22, YLMYY22, CHH^{+}24, Don20, IW20, MSF^{+}22, NNS^{+}22, ON21, Pol23, RP23, ZX20, WZ23]. evidence-based [NNS^{+}22]. Evidence-neglect [Mar23]. Evidences [NCTd20]. Evolution [GRS24, XLL22, AS20]. examination [BBAK21]. Examining [Abd23]. excellence [GWS23, SFvD^{+}21, Ter23]. exist [Vin22]. expansive [VSGMF21]. expectations [ALK22, FvL22]. experience [DSW23]. Experiences [AGKL21, HPV^{+}22, NCTd20, RO24, Ter23]. experiential [SU23]. experiment [WZ23]. experimentation [CLiM^{+}24, GKR^{+}21, RKT23]. expert [WvSG22]. Experts [Loy22, YXX21]. exploration [RAB^{+}23]. explore [Cha22]. Exploring [HPV^{+}22, Lil20, PGM23, RR20, WDJ^{+}21, vdKYYWR23, AOF22]. external [AOF22, IW20].

facet [Her23]. facilitate [CCT21a, CCT21b]. Factors [Rea22, CBUS23, NNS^{+}22]. faculty [AD20]. failure [LSvH22]. fair [HTB22]. fairness [AOF22]. farmer [ZYH^{+}20]. favorable [KH22]. favoritism [AD20]. FDI [YH21]. fears [Kim21]. federal [GRL23]. federalism [ZFCL21]. Federally [Lin21]. fertilising [NJV23]. few [SFvD^{+}21]. field [HSSS21]. fields [dlTGRS21]. Fifty [BSTG21]. finance [Her23]. Find [Lag21]. Finland [AS20, KLL23, SC21a]. Finnish [PEG20]. Firm [AMDN21, FS20, dSTRC20, ZW20]. firms [GSR22, THPT20]. fiscal [ZFCL21]. fisheries [OdS20]. Flagship [MY21]. Flemish [PEG20]. flexible [XG23]. flow [dRS20]. flows [NCTd20]. focus [Kis20, LER^{+}22]. Follow [Boz22]. food [AGKL21]. Force [CBUS23]. Foreign [GSR22]. formal [THPT20, WS23]. formation [MGM^{+}21, PV24, vWBT20]. Forward [NK21, JTWW21]. foster [AD20]. Fostering [DC23, CLiM^{+}24]. Foundation [OF22, PV24]. four [Tas24]. four-asset [Tas24]. frames [CCF24]. Framework [PBSA20, AY23, CP23, GFF^{+}23, LKZC22, MA22, SMPB24, WWF^{+}20, dlGGM21]. frameworks [Bar23]. framing [RK23]. France [SWAD24]. freedom [FKF^{+}23]. fresh [LCL23]. friction [PBW20].

Frictions [Sol21]. Fun [SU23]. Functional [EZI20]. functioning [MGM^{+}21]. Funded [Lin21, BL22, LW21, MBT20, OF22, WB22]. Funding [APV21, SFvD^{+}21, Abd23, Bar23, FS20, GRL23, GST^{+}20, IW20, LM23, Phi22, PBSA20, PEG20, SK23, SU23, SFP23, ZXW^{+}20, Zou22]. Funds [SC21a]. future [Kim21, RAJV21]. futures [ALK22, JL21, Law24, RP23, RR20]. Fuzzy [AVZ21].

gap [LLRD23, SFP23]. GDPR [MRB20]. Gender [Lag21, PV24, SK23, CPT^{+}24, PH23, SPB23, SFP23, VV22]. Gene [NK21, SJ24, WBM22]. Gene-Edited [NK21]. general [RK23]. generalizability [SX19, SX20]. generating [dJBN22]. generational [Sil24]. genetically [Pan20]. Genome [SWA^{+}21, WS21]. German [SU22]. Germany [CHH^{+}24, FLBM23, RR20]. Gerontocracy [Sil24]. Ghana [AYDQ^{+}21]. Global [YXX21, BSTG21, BMA23, Qia23, KR23, RK23, RRC^{+}23]. globalization [RAB^{+}23]. Globally [MY21]. goals [BMA23]. goods [WS21]. govern [CP23]. Governance [HL20, Yan21, YXX21, Abd23, AGKL21, CP23, GRS24, OSV^{+}23, PSK^{+}23, WLW20]. governed [CP23]. Governing [WHL^{+}21]. Government [Qia23, BL22, BZG21, DSW23, Pen23, ZX20, ZFCL21]. government-funded [BL22]. governmental [Dui20]. governs [OSV^{+}23]. graduate [CH20a, CH20b]. grand [BS22b, HL20]. grant [HTB22, LRS24, LMS21, MA22, SU23]. grants [AMD24, PV24]. greater [ASFB^{+}23, SC22]. green [TSS23, TTF20]. group [LER^{+}22]. groups [SFvD^{+}21]. growing [GRS24]. growth [HJMX23, HYO^{+}21a, HYO^{+}21b, SWA^{+}21, Tas24]. Guangdong [CZtB21, THPT20]. Guidance [HLL^{+}23]. hailing [SX19, SX20]. haven [YH21]. Heads [Lag21]. health [BBAK21, HPV^{+}22, Mar23, SPB20, SU22, TRdRABS22, WS21, vdKYYWR23]. healthcare [MGM^{+}21, MAH^{+}20]. Heart [NLOL24]. helix [PBW20, CCT21a, CCT21b]. help [SHK^{+}20]. Herding [GZG^{+}23]. heterogeneity [Ano23c, VBFLP22]. hidden [FZW22]. Hierarchy [Yeh21, Lin20]. High [SWA^{+}21, Yeh21, LXLY20, Pen23]. High-Energy [Yeh21]. High-tech [SWA^{+}21, LXLY20, Pen23]. higher [ABM21, Her23]. historical [RAB^{+}23]. HIV [XHQ^{+}22]. HIV/AIDS [XHQ^{+}22]. Home [MY21]. homegrown [Pen23]. honest [Bav23, SHK^{+}20]. Hong [SC22]. Hopes [ALK22]. Horizon [EG20]. hotspots [Cha22]. house [THPT20]. hubs [Hem23]. human [BPS20, GCGAMASG21, ISS22, WS21]. hybridisation [NJV23]. Hypergraphs [BMCF21]. hypothesis [YH21].

I-Corps [YKW21]. ICT [EG20, KCL20, Kim21]. idea [AS20]. identification [PEG20]. identity [hKZ22]. Ideology [HV23]. idiosyncrasies [dSTRC20]. IIB [NLOL24]. Imaginaries [Pes21, Rud22]. imaginative [Law24]. imagined [HSAR23]. IMI [GFF^{+}23]. IMI-PARADIGM [GFF^{+}23]. Immediacy [WB22]. IMPA [PDM23]. Impact [AMDN21, CM21, TBM21, Yan21, YKW21, AMD24, ABS21, ASFB^{+}23, BMCF21, CH20a, CH20b, HR20, Lau20, LMS21, MA22, Rea22, dJB22, dJBN22]. Impacts [WB22, Bre22, OF22, SFQdM24, XHQ^{+}22]. implement [KKS20]. implementation [BPS20, CM20, Lil20, WLW20]. Implications [MRB20, RAM22, Mou24, PH23, XLL22]. importance [Rea22]. improve [ZX20]. improved [MAH^{+}20]. Improving [TS23]. in-house [THPT20]. inbreeding [MSF^{+}22]. incentive [LL22]. incentives [Ehs21]. Inclusion [WB22, LMS21]. Inclusion-Immediacy [WB22]. incoherent [WHH24]. incomplete [RCK23]. Increasing [BFH^{+}20]. India [SJ22]. indicators [Mou24, PGM23]. indirect [GRFBL21]. Individual [MG22]. Indonesia [NNS^{+}22]. industrial [RKT23, THPT20, TSS23]. industries [Pen23]. Industry [AMDN21, Kim21, BSTG21, FMF20, HJMX23, LC23, LLM^{+}22, MVNSdUC22, NNAH22, SDM23, VJMHdLNA20, WJZ20, dSFSN23]. inefficiencies [FD24]. inequalities [WB22]. Influence [Yan21, GV20, Kis20, dJB22]. Influences [YXX21]. informal [SJ22]. information [Pan20]. Infrastructure [FKS^{+}21, hKZ22, KKS20, SWA^{+}21]. infrastructures [OI22]. initiative [SC22, DF22, SWA^{+}21]. initiatives [DWK23]. Innovating [MAH^{+}20]. Innovation [AVZ21, LSvH22, PL22, TS23, vWBT20, AMD24, AYDQ^{+}21, AS20, AKC22, Ano23b, AFF20, AIIOL22, Ben22, BZG21, BS22b, BS21, CM20, CCT21a, CCT21b, CK21, CNA^{+}20b, CNA^{+}20a, CNA^{+}20c, Clo21, Coh22, CLiM^{+}24, DSW23, DWK23, DS23, EOS24, EZI20, Ehs21, FZW22, GSR22, GRS24, GKR^{+}21, Hem23, Hen23, HSYW21, HJMX23, JTWW21, JWFY22, KJR20, LLM^{+}22, LBCG^{+}23, LXLY20, LKZC22, LL22, MGM^{+}21, MAH^{+}20, MRB20, NLP23, Nik24, NJV23, ORZJU24, Pes21, PFPV20, RKT23, RCK23, SX19, SX20, SC22, SJ22, SGDW23, dMSVF23, Sol21, SMPB24, THPT20, VV22, WJZ20, WWF^{+}20, WHL^{+}21, WZ23, XLL22, XG23, YLGZ23, YAB^{+}20, ZX20, ZFCL21, ZYH^{+}20, ZW20, DF22, SBH^{+}22]. innovators [ST21]. inputs [LCL23]. Insights [HHL21, LER^{+}22, SC21a, YAB^{+}20, BFH^{+}20]. Institute [NLOL24]. institutes [CZtB21, Kru20]. Institution [THPT20, ABM21]. Institution-driven [THPT20]. Institutional [LC23, ASFB^{+}23, CCT21a, CCT21b, Coh22, CLiM^{+}24, GST^{+}20, HL22, MSF^{+}22, MP23, Rea22, SMPB24, ZW20, SFQdM24]. institutionalist [Ben22]. Institutionalizing [Coh22]. Institutionally [Dui20]. institutions [Abd23, DC23, PDM23]. instruments [BS22b, CZL^{+}20, LL22, dSFSN23]. insurance [WZ23]. integrated [CP23]. integration [BCFU23, PM20]. integrity [DC23, HLL^{+}23, LER^{+}22]. Intellectual [ZW20, Par20, Sid20]. intelligence [AY23, LLM^{+}22, OSV^{+}23, PM20]. intensity [MPCAC20]. Inter [ZYH^{+}20, HL20]. inter-organizational [HL20]. Inter-provincial [ZYH^{+}20]. interact [ZG22]. interaction [dJBN22]. interactions [TG22]. interactive [SJ22]. interdisciplinarity [Ano23c, hKZ22, VBFLP22]. interdisciplinary [KKS20, hKZ22]. interest [Bav23]. Interests [Yeh21, CK21, FKF^{+}23]. interface [BTT22, OdS20]. intergovernmental [RRC^{+}23]. intermediaries [NLP23, SDM23, vWBT20]. International [DP21, ASFB^{+}23, Cha22, EKFR^{+}23, HHL21, vWBT20]. internationalization [MSF^{+}22]. Interpreting [HÅ23]. Interreg [MP23]. interventions [PSK^{+}23]. introduction [KR23]. invasive [Haw20]. Invention [LS20]. inventions [NNAH22]. Inventor [HSYW21]. inventors [MG22]. investigation [Woe21, ZX20, ZQH^{+}23]. Investment [AFF20, HYO^{+}21a, HYO^{+}21b, JWFY22]. investments [HR20]. invisible [Rud22]. involvement [Ano23b, HPV^{+}22, LBCG^{+}23, SPB20]. issue [Pes21, SHK^{+}20]. issues [ISS22]. Italian [AD20, NPZ21]. Italy [ABS21].

Jakarta [NNS^{+}22]. Japan

[IW20, NNAH22, TO21]. Japanese

[ON21, OI22]. journal [WHH24]. journals [LM23]. Judgement [YXX21].

Key [RR20, KKK^{+}21, YCL22, Woe21]. Key-Narratives [RR20]. know [CL20]. Knowledge [Ano22, FE20, Lin21, OdS20, TRdRABS22, Yan21, wY22, dlTGRS21, APGL20, AGKL21, Ano23d, CZtB21, FMF20, HV23, HL21, JGK22, LS20, MCDD23, NCTd20, VV22, VBK^{+}23, ZS21, dRS20]. knowledge-based [CZtB21]. knowledge-brokering [Ano23d, MCDD23]. Kong [SC22]. Korea [KKK^{+}21, KCL20]. Korean [BL22, CK21, Mo23]. Kuwait [AIIOL22].

labor [NRSA23, Sil24]. laboratories

[YCL22]. Laden [YXX21]. Lamar [OF22]. landscape [AKC22]. large [KKS20]. large-scale [KKS20]. Latin [EKFR^{+}23, Ter23]. Layered [RCK23]. leap [JWFY22]. learned [HLL^{+}23]. learning [Don20, RKT23, SJ22, SC21b]. leaving [NPZ21]. Legal [SMPB24]. Legislative [AAW^{+}20]. legitimacy [DWK23, WvSG22]. Legitimate [RML^{+}21, SJ24, VPK20]. lens [Cav24, PH23]. less [SU23]. Lessons [Boz22, HL20, OdS20, Ehs21, MAH^{+}20]. library [SGW21]. Lines [NK21]. Listing [WHH24]. lists [WHH24]. literature [GRS24, KJR20, MVNSdUC22]. Local [JWFY22, Pen23, TO21, Yan21, ZFCL21, BMA23, HSAR23, ON21, SGDW23]. localization [HL21]. logics [LC23]. long [RO24, YLGZ23]. long-range [YLGZ23]. long-term [RO24]. lottery [HTB22, Phi22]. low [VJMHdLNA20]. Lung [NLOL24].

made [CM21]. Madrid [QS23]. mainstreaming [SK23]. make [HTB22, JS23, LMS21, SJ24]. Maker [FvL22]. makes [SFQdM24]. Making [APGL20, Ano23d, BBAK21, FKS^{+}21, HÅ23, MCDD23, NNAH22, SJ24, Wil22, HLM^{+}23, LKZC22, SPB23, dJB22]. management [AYDQ^{+}21, BPS20, IW20]. managerial [HLS20]. mandates [Kru20]. manufacturing [NLP23, SJ22]. mapping [LM23]. Market [AVZ21, BWG22, BGW22, NLOL24, Pen23, Sil24]. marketization [LL22]. markets [Sid20]. matching [MG22]. Materials [SWA^{+}21]. matter [Don20]. matters [SGW21]. Matthew [ZXW^{+}20]. meaning [Pra20]. Meaningful [Lag21, HPV^{+}22]. measurement [GCGAMASG21]. mechanisms [HL20, LM23]. Media [MB20]. medical [KM23]. medicine [BBAK21, DS23, Sid20, Pra20]. medium [Ehs21]. medium-sized [Ehs21]. meetings [PH23]. Meets [ZAB^{+}20, ZAB^{+}21]. mental [vdKYYWR23]. merrier [HSYW21]. Meta [AMDN21]. Meta-Regression [AMDN21]. Methods [AAW^{+}20]. metrics [Wil22]. micro [Nel24]. micro-dynamics [Nel24]. Microalgae [RR20]. mid [LM23]. mid-size [LM23]. migration [NPZ21]. Mission [AVZ21, NJV23, RCK23, BMA23, BS21, JTWW21, hPKvR22, SC21a, SMPB24, WWF^{+}20, WHL^{+}21]. Mission-Oriented [AVZ21, NJV23, BMA23, BS21, JTWW21, SMPB24, WWF^{+}20, WHL^{+}21]. Missions [JWT^{+}23, RO24]. mix [AMD24, CL20, Hen23, KLL23, ON21, SGDW23, WJZ20]. mixes [DSW23]. mobility [CK21, MSF^{+}22]. model [BS22a, CCT21a, CCT21b]. modelling [BCFU23]. models [CZtB21]. Moderating [LL22, THPT20, BZG21, Qia23]. moderators [SFQdM24, dMSVF23]. modern [Sil24]. modes [HHL21]. modified [Pan20]. monitoring [Mou24, PGM23, RCK23]. monoclonal [dSFSN23]. motivates [AOF22]. motivation [SU22]. motivational [AOF22].

Motivations [DP21, Nel24]. move [FMF20]. movement [RP23, FvL22]. Moving [Mou24]. Mozambique [ZS21]. Multilevel [AMD24, IW20, SGDW23, ON21]. multiple [HL22, Kru20]. Myriad [Gho20].

name [CNA^{+}20b, CNA^{+}20a, CNA^{+}20c]. nanoELSI [hKZ22]. Narratives [RR20, RP23]. narrowing [KM23]. Nation [SLSC^{+}21, RK23]. nation-state [RK23]. Nation-Wide [SLSC^{+}21]. National [CM20, OF22, WDJ^{+}21, AMD24, ABS21, AS20, NCTd20, NJV23, SGDW23, ZX20, NLOL24]. nations [dRS20]. natural [ZXW^{+}20]. Navigating [RO24]. nearly [NRSA23]. need [vdKYYWR23]. needs [HPV^{+}22]. neglect [Mar23]. Negotiating [Clo21, VBK^{+}23]. nesting [BS22b]. Network [TBM21, Cha22, MG22, TO21, WLW20, XLL22]. networks [Bre22, CCT21a, CCT21b, EG20, PGM23, RKT23, TRdRABS22]. NHS [MAH^{+}20]. non [Haw20, LMS21]. non-academic [LMS21]. non-invasive [Haw20]. Nordic [SBH^{+}22]. Normal [RML^{+}21]. normative [VSGMF21]. North [Bar23]. Novel [APV21]. NSF [WB22]. NSF-funded [WB22]. numbers [JS23]. Numeric [JS23]. nurturing [NLP23]. Nutrition [RR20].

object [Ano23a, EET22]. objects

[JWT^{+}23, SHK^{+}20]. Observable [SFP23]. observing [CBUS23]. occurrence [NNS^{+}22]. Ocean [Pol23]. off [RKT23]. offensive [WZ23]. offs [HYO^{+}21a, HYO^{+}21b]. Online [PH23, REL22, Rzy22, VPK20]. Open [MA23, ABC^{+}21, BS22a, HJMX23, LC23, MRB20, NN22, VSGMF21, dlGGM21, Ano23e, Lil20]. Opening [VSGMF21, BCFU23]. Openness [DS23, Sid20, Pra20]. opportunities [AJG22]. Opportunity [RMP23]. options [Bla21b, Bla21a, TTF20]. Orchestrating [EG20]. orderings [Law24]. organisation [SCY21]. Organisations [Bor21, Ano23d, HHL21, LKZC22, MCDD23, dlTGRS21]. Organizational [FKF^{+}23, dSTRC20, AOF22, BZG21, HL20, dJB22]. organizations [CPT^{+}24, PDM23, RRC^{+}23, dJBN22]. Oriented [AVZ21, Bor21, BMA23, BS21, JTWW21, NJV23, SMPB24, WWF^{+}20, WHL^{+}21]. other [TG22, Ter23, YLMYY22, ZG22]. outcomes [GKR^{+}21, PFPV20]. Overcoming [CK21]. Overlapping [KCL20]. overview [BSTG21, Rud22]. owned [GSR22].

pan [HLL^{+}23]. pan-European [HLL^{+}23]. pandemic [Boz22, FLBM23, WvSG22]. panel [PH23]. panels [ASGRM23, LMS21]. PARADIGM [GFF^{+}23]. paradox [FD24]. Parliament [ZAB^{+}20, ZAB^{+}21]. Partial [HTB22]. participation [Cav24, VPK20]. Particle [Yeh21]. partnership [BTT22]. past [JL21, Kim21]. Patent [NCTd20, WZ23]. Patents [Haw20, Cha22, MG22, Sid20, XLL22, ZQH^{+}23]. Paths [NK21]. pathways [WWF^{+}20]. Patient [SPB20, GFF^{+}23, HPV^{+}22]. Patterns [APV21, NPZ21]. Pay [APV21]. Peer [YCL22, ASGRM23, BS22a, LRS24, PH23, Phi22, PEG20]. peer-reviewed [PEG20]. perception [SWAD24]. Perceptions [CNA^{+}20b, CNA^{+}20a, CNA^{+}20c]. perform [SHK^{+}20]. Performance [AMDN21, WDJ^{+}21, Yan21, JWFY22, KH22, PEG20, WLW20, Zou22]. performance-based [PEG20]. performing [CPT^{+}24]. period [NCTd20]. peripheral [SC21a]. Personalised [Pra20]. personalized [BBAK21]. Perspective [BTT22, Bav23, LMR^{+}20, Ben22, BZG21, BS22b, GV20, RK23, SPB20, TTF20, ZFCL21]. Perspectives [AYDQ^{+}21, Hem23, KM23, FLBM23, MAH^{+}20]. Peru [NRSA23]. Phase [NLOL24]. Physics [Yeh21]. pick [DSW23]. plan [YLGZ23]. platforms [MGM^{+}21, REL22]. play [Bav23]. plural [Ano23c, VBFLP22]. Policies [Lag21, Ano23e, CBUS23, CCF24, GV20, Hen23, JLX21, LER^{+}22, LCL23, MA23, SX19, SX20, SC22, SCY21, VPK20, WHL^{+}21, ZG22, dSFSN23]. Policy [Abd23, AAW^{+}20, AVZ21, Ano23a, MY21, MP23, PL22, SPB23, TTF20, AD20, AMD24, APGL20, AYDQ^{+}21, AS20, Ano23d, AFF20, BTT22, BCFU23, Ben22, BPS20, BS22b, Boz22, BMA23, CM20, CZL^{+}20, CL20, DSW23, DF22, Don20, DS23, EOS24, EET22, FvL22, GKR^{+}21, GZG^{+}23, HV23, Hen23, HÅ23, HLM^{+}23, HYO^{+}21a, HYO^{+}21b, JTWW21, JWT^{+}23, JS23, Kis20, KLL23, LLM^{+}22, LLRD23, Lil20, LXLY20, LMR^{+}20, MCDD23, Mar23, MB20, Mo23, Nik24, NJV23, ON21, OdS20, RKT23, RCK23, RMP23, Rzy22, SPB20, SBH^{+}22, SGDW23, Sol21, Tas24, TS23, TG22, TSS23, WJZ20, WWF^{+}20, WLW20, XLL22, YLGZ23, dlGGM21]. Policy-making [SPB23, HLM^{+}23]. policy-nesting [BS22b]. policymakers [ZAB^{+}20, ZAB^{+}21]. political [Ben22, CK21, FZW22, JWFY22, Sid20]. politicisation [Loy22]. politics [Law24, WvSG22]. pollution [YH21]. portfolio [PGM23]. portfolios [vdKYYWR23]. position [Qia23]. positioning [ABM21]. positivity [NNS^{+}22]. Post [RML^{+}21, Gho20]. post-Myriad [Gho20]. Post-Normal [RML^{+}21]. potentials [SBH^{+}22]. power [SPB23, WJZ20]. practice [AGKL21, BFH^{+}20, BMA23, hKZ22, Lil20]. Practices [Yeh21, ABC^{+}21, BTT22, CNA^{+}20b, CNA^{+}20a, CNA^{+}20c, Dui20, Fli22, LC23, RCK23, Sol21, dJB22]. practitioner [Ano22, wY22]. practitioners [SHRA20]. precarity [FD24]. precarity-driven [FD24]. Precision [Pra20]. preferences [Vin22]. premises [JTWW21]. prenatal [Haw20]. presents [Law24]. pressure [CH20a, CH20b]. Priorities [vdKYYWR23, ST21]. Prioritizing [GRL23]. priority [AFF20]. private [GSR22, Qia23]. probabilities [PBSA20]. problem [WWF^{+}20]. process [BZG21, GST^{+}20, NJV23, SPB23, dMSVF23]. processes [FMF20, PFPV20, SPB20]. processual [PBW20]. procurement [EZI20, KJR20, LKZC22]. production [AGKL21, MBT20, OdS20, TTF20, VV22, VBK^{+}23]. Productive [TG22, dJBN22]. productivity [Mo23]. professional [BL22, ZYH^{+}20]. professionalism [BS21]. professionals [IW20]. profiles [dlTGRS21]. program [LSvH22, XHQ^{+}22]. programmatic [AS20]. programme [RO24]. programmes [SPB23, SGDW23]. programs [HL20, PBSA20]. progress [KM23]. project [BL22, Bar23, Cav24, GFF^{+}23, LSvH22, Nel24]. projects [BL22, Bar23, EG20, NLOL24, Vin22]. promise [AY23, FvL22, WLW20]. promises [BBAK21, JTWW21, JL21]. promising [RAJV21]. Promissory [WS21]. promote [LXLY20, WZ23, XG23]. Promoting [Kim21]. promotion [CNA^{+}20b, CNA^{+}20a, CNA^{+}20c]. property [Par20, Sid20, ZW20]. proposal [GCGAMASG21]. proposals [ASGRM23, PBSA20, SK23]. prosocial [SU22]. Prospect [Sid20]. Protection [Loy22]. provided [HLL^{+}23]. provincial [ZYH^{+}20]. Public [Bor21, KJR20, SWAD24, WBM22, dSFSN23, Abd23, AY23, Ano23b, AFF20, Boz22, CP23, CLvH22, Coh22, CLiM^{+}24, EOS24, FS20, FD24, Kru20, LM23, LBCG^{+}23, LKZC22, LMR^{+}20, Nel24, ON21, Pan20, Pes21, RR20, SCY21, SC21b, WvSG22, WS21, ZQH^{+}23, dlTGRS21]. publications [ABS21, PEG20]. publicly [LW21]. publicness [LW21]. publics [WS21]. publishers [LM23]. publishing [RAM22]. pulse [Fli22]. push [HL21].

quadruple [PBW20]. qualitative [Par20]. quality [FKF^{+}23, HSYW21, WHH24]. Quantitative [BFH^{+}20]. quasi [WZ23]. quasi-experiment [WZ23]. Questioning [RAJV21].

R&D [AMD24, AIIOL22, ASFB^{+}23, BL22, BWG22, BGW22, CZtB21, FS20, GRL23, HL20, HYO^{+}21a, HYO^{+}21b, JWFY22, KH22, Mo23, MPCAC20, ON21, Qia23, SCY21, TO21, THPT20, Zou22]. R&I [RO24]. randomly [Phi22]. range [YLGZ23]. rationales [GV20]. re [FE20]. re-entanglement [FE20]. reading [HSSS21]. realize [HPV^{+}22]. really [SCY21]. reasons [LRS24]. recovery [AFF20, GRFBL21, KLL23]. recruitment [AD20]. REF [BMCF21]. reflection [JTWW21]. reflexivity [GST^{+}20]. reforms [AD20]. regimes [FE20, RAJV21, WHH24, WS21]. region [VJMHdLNA20, QS23]. regional [AMD24, Ben22, CZtB21, Hen23, TO21, TTF20, VV22, YLGZ23, ZFCL21, TS23]. regions [BMA23, SC21a]. Regression [AMDN21]. Regulation [NK21, ISS22, MB20, WBM22]. regulations [RAM22, YH21]. regulatory [Hem23]. Rejecting [SC21b]. related [Ano23e, JLX21, MA23]. relations [SGW21, TCWZ21]. Relationships [AMDN21, VJMHdLNA20, ZAB^{+}20, ZAB^{+}21]. relevance [BFH^{+}20]. reluctant [BWMH22]. remedies [LRS24]. renewable [KH22]. reorienting [GKR^{+}21]. Republic [GV20]. requirements [REL22]. Research [AAW^{+}20, Bor21, CM21, DP21, FKS^{+}21, GWS23, Lin21, LSvH22, Phi22, SLSC^{+}21, TBM21, XHQ^{+}22, Abd23, ASGRM23, AJG22, ABS21, BTT22, Bar23, BMCF21, BS22b, Bre22, CPT^{+}24, Cav24, CLvH22, CNA^{+}20b, CNA^{+}20a, CNA^{+}20c, CL20, Coh22, CLiM^{+}24, DC23, Dui20, FD24, GRFBL21, HPV^{+}22, HLL^{+}23, HHL21, IW20, JWT^{+}23, KKK^{+}21, KKS20, KM23, Kis20, Kru20, hKZ22, LER^{+}22, LC23, Lau20, Lil20, LW21, Nel24, Nik24, NNAH22, OI22, OF22, Par20, PEG20, Rea22, REL22, SFvD^{+}21, SPB20, SK23, SCY21, SU23, VV22, Wil22, YLGZ23, YLMYY22, Zou22, dJBN22, dlTGRS21, vdKYYWR23]. research-performing [CPT^{+}24]. Researcher [MSF^{+}22, PSK^{+}23, Bar23, HPV^{+}22, HSAR23, NPZ21]. Researchers [SLSC^{+}21, CHH^{+}24, HPV^{+}22, SFP23]. resilience [KLL23]. resistance [Abd23, BWG22, BGW22, SPB23]. Resource [Lin20, BPS20]. resources [YLMYY22]. Responding [FLBM23, YLMYY22]. response [MB20]. Responses [RML^{+}21, GZG^{+}23]. responsibility [GST^{+}20, HHL21, Nik24, RMP23]. responsible [CNA^{+}20b, CNA^{+}20a, CNA^{+}20c, PM20]. Restoring [BGW22, BWG22]. Restricted [SX19, SX20]. results [MPCAC20, SLSC^{+}21]. rethink [Rzy22]. Retraction [Rzy22]. returnee [YCL22]. Returning [CBWJ20, MY21]. Reverse [MY21]. review [ASGRM23, BS22a, CH20a, CH20b, HLL^{+}23, KJR20, LRS24, Par20, PH23, Phi22]. reviewed [PEG20]. reviewers [LMS21]. revisited [YH21]. revolution [TSS23]. rights [Sid20]. rise [AJG22]. risk [BZG21, Nik24]. Role [BL22, APGL20, ABM21, ABC^{+}21, Bav23, GRS24, NLP23, ORZJU24, PL22, Qia23, Vin22, dJBN22, dlTGRS21, vWBT20]. roles [Ano23b, HLS20, LBCG^{+}23, PSK^{+}23, SHK^{+}20].

S&T [SFP23]. Sand [NK21]. sandbox [Hem23]. sanitation [vWBT20]. SBIR [LSvH22, NLOL24]. scale [ABC^{+}21, DSW23, GCGAMASG21, KKS20]. scale-up [DSW23]. scanning [BZG21]. scheme [CPT^{+}24, CM20, Mo23]. scholarly [LM23, RAM22]. scholars [MBT20]. scholarship [Par20]. Science [AAW^{+}20, Ano23a, APV21, BPS20, DP21, KR23, LMR^{+}20, OF22, PV24, RP23, RK23, RRC^{+}23, WvSG22, WLW20, WDJ^{+}21, ZAB^{+}20, ZAB^{+}21, Ano23e, ABC^{+}21, Bav23, BFH^{+}20, Bla21b, Bla21a, Boz22, CBWJ20, Cav24, CM20, Don20, DS23, EKFR^{+}23, EET22, FKF^{+}23, Fli22, HV23, HÅ23, HL21, HLM^{+}23, LC23, Law24, LLRD23, MA23, NN22, OdS20, Pol23, Rea22, RAJV21, RAB^{+}23, SHK^{+}20, Sil24, SU22, TCWZ21, VSGMF21, WS23, ZXW^{+}20, Lil20, Vin22]. science-related [Ano23e, MA23]. Sciences [CM21, TRdRABS22]. Scientific [EKFR^{+}23, WDJ^{+}21, GWS23, JL21, Kis20, LS20, MBT20, Nel24, PDM23, RP23, SWAD24]. scientist [BS22a, Phi22, SJ24]. Scientists [Yan21, APGL20, CBWJ20, OdS20, YCL22, ZAB^{+}20, ZAB^{+}21]. scoping [HLL^{+}23]. Scoreboard [TS23]. search [BZG21]. sector [Ano23b, CP23, Clo21, LBCG^{+}23, NLP23, PM20, RO24]. seduction [Abd23]. seen [TG22]. selection [Woe21, ZXW^{+}20]. Selectivity [APV21]. self [SC21b]. self-driving [SC21b]. Semantic [BMCF21]. semiconductor [HJMX23]. sense [APGL20, Ano23d, HÅ23, HSAR23, MCDD23, Wil22, dJB22]. sense-making [dJB22]. servants [BWMH22]. Service [FKS^{+}21, LCL23]. services [OI22, VPK20]. setting [SPB20]. Shall [Rzy22]. shape [LKZC22]. shapes [DSW23]. shared [GFF^{+}23]. sharing [MRB20]. Shenzhen [SC22, Zou22]. shift [NN22]. Shipbuilding [AVZ21]. side [TG22]. simulation [BS22a]. Singapore [HL21, NLP23]. Situated [VV22]. size [HSYW21, LM23]. sized [Ehs21]. slowdown [LS20]. Small [LSvH22, Ehs21, LM23]. small- [Ehs21, LM23]. smart [Ben22, Mou24]. SME [Hen23]. SMEs [ON21, XG23]. Smith [OF22]. Social [EOS24, DWK23, Rea22, Vin22]. societal [Bre22, Lau20, ZS21, dJB22, dJBN22]. societies [HLL^{+}23]. society [JGK22, Law24, SU22]. socio [RP23]. socio-ecological [RP23]. sociologists [ABS21]. Sociology [DP21]. Sociotechnical [MAH^{+}20, PGM23, Rud22]. solution [RKT23, WWF^{+}20]. solutions [VPK20]. Sources [WDJ^{+}21, dRS20, KH22]. South [AGKL21, Bar23, KR23, KCL20, Kru20, RK23, RRC^{+}23]. Southeastern [YKW21]. space [Clo21, SJ24, VBK^{+}23, WWF^{+}20]. Spain [QS23]. spatial [CLvH22]. speak [HSSS21]. specialisation [dlTGRS21]. specialization [Ben22, Mou24]. specific [HLL^{+}23]. stability [HYO^{+}21a, HYO^{+}21b]. stakeholder [JWT^{+}23]. stakeholders [KKK^{+}21, RAM22]. stand [PFPV20]. State [Bor21, GSR22, OI22, RK23, YCL22]. state-of-the-art [OI22]. state-owned [GSR22]. statements [BMCF21]. Staying [NPZ21]. STEM [SK23]. STI [GV20]. STIP [GRS24]. story [JGK22]. strategic [ABM21, BPS20]. Strategies [FKS^{+}21, Bar23, Bre22, FKF^{+}23, MP23, ORZJU24, dJB22]. strategy [DF22, Pen23]. strengthen [Ano23c, VBFLP22]. Strengthening [hPKvR22]. stress [CH20a, CH20b]. Structural [SC21a, WDJ^{+}21, CM20, CLiM^{+}24]. structure [Cha22, VSGMF21]. students [CH20a, CH20b]. studies [HSSS21]. Study [DP21, TBM21, AIIOL22, BTT22, CK21, FZW22, Kru20, LER^{+}22, LLRD23, LSvH22, Loy22, MG22, Mo23, NLP23, SK23, Sol21, ZS21, ZQH^{+}23, Zou22]. style [ZFCL21]. subsidiaries [GSR22]. subsidies [ON21, Qia23]. Subsidy [LCL23, BZG21]. success [MVNSdUC22]. successful [AD20]. Superclusters [DF22]. supply [LCL23]. support [CLvH22, LLM^{+}22, Pen23, TO21]. Supporting [HPV^{+}22, MGM^{+}21]. Survey [SLSC^{+}21, Phi22, RR20]. sustainability [HSSS21, Mou24]. sustainable [VSGMF21]. Swedish [BMA23, RCK23, SGW21]. Switching [Clo21]. Switzerland [BBAK21]. system [AS20, CBWJ20, LLM^{+}22, LCL23, NJV23, PGM23, vWBT20]. systematic [CH20a, CH20b, KJR20]. Systems [AIIOL22, MAH^{+}20, PEG20].

Tacit [HL21]. Taiwan [CCT21a, CCT21b]. Taking [Fli22]. Talents [MY21]. talk [SBH^{+}22]. task [Qia23]. tasks [BWMH22]. tax [Ehs21]. taxonomy [SDM23]. team [HSYW21, PV24]. tech [LXLY20, Pen23, SWA^{+}21]. technical [BPS20, KCL20, ZYH^{+}20]. techno [JL21]. techno-scientific [JL21]. Technological [Cha22, ZX20]. Technology [Bor21, CZtB21, Hem23, Bla21b, Bla21a, CM20, Cha22, Hen23, HLM^{+}23, JLX21, NLOL24, NRSA23, Pan20, Pes21, PM20, RAJV21, SCY21, SJ24, Tas24, WS23, WLW20, dSFSN23]. technology-based [Tas24]. term [RO24]. testing [Haw20]. Their [Yan21, YXX21]. therapeutic [dSFSN23]. there [FZW22, Haw20, NRSA23]. these [Ter23]. third [hPKvR22, SC21a]. those [ASFB^{+}23]. Threat [Lil20, KM23]. three [BMA23]. threshold [CZL^{+}20]. TMNs [Ano23a, EET22]. together [APGL20]. tool [WS23]. top [MBT20]. top-funded [MBT20]. traceability [LCL23]. Tracing [RMP23]. trade [HYO^{+}21a, HYO^{+}21b, RKT23]. trade-off [RKT23]. trade-offs [HYO^{+}21a, HYO^{+}21b]. Trading [Mo23]. training [LER^{+}22]. trajectories [KM23]. transdisciplinarity [Ano23c, HÅ23, VBFLP22]. Transdisciplinary [AAW^{+}20, AGKL21]. transfer [Ano22, Cha22, CZtB21, FMF20, SCY21, wY22, dlTGRS21]. Transfers [Lin21]. Transform [FKS^{+}21, RO24]. transformation [CCF24, EOS24, SPB23]. Transformative [AYDQ^{+}21, GKR^{+}21, PL22, BS22b, JWT^{+}23, KLL23, RCK23, SBH^{+}22]. transforming [NJV23]. transition [BWMH22]. transitions [CHH^{+}24, PGM23]. translation [SHRA20, SJ24]. translations [Sol21]. transparency [BS22a, MRB20, NN22, RAM22]. transversal [HSSS21]. trap [AIIOL22]. Trends [AJG22]. trial [MRB20, RAM22]. Triangle [Nik24]. Triggering [NK21]. Triple [CCT21a, CCT21b]. Triple-Helix [CCT21a]. turbulence [THPT20]. Turkey [Ano23a, EET22]. Turning [Pes21]. turnover [JWFY22]. two [Don20]. typology [WHL^{+}21].

UK [AJG22, Ehs21, Mar23]. Unboxing [JGK22]. uncertainties [Lau20]. uncertainty [FLBM23, JGK22]. undermine [BS22a]. Understanding [Ano23b, BWMH22, CHH^{+}24, JWT^{+}23, LBCG^{+}23]. understandings [Ano23c, VBFLP22]. Unite [RP23]. universal [Hem23]. Universities [Cav24, CM21, PL22, CHH^{+}24, DWK23, GCGAMASG21, HLS20, IW20, KKK^{+}21, Lin20, Woe21, dJB22]. University [AMDN21, BSTG21, SDM23, YKW21, CCT21a, CCT21b, FMF20, FD24, LC23, MVNSdUC22, NNAH22, hPKvR22, QS23, SC21a, SFQdM24, VJMHdLNA20, ZQH^{+}23, PBSA20]. university-centric [CCT21a, CCT21b]. unleash [SBH^{+}22]. unobservable [SFP23]. upgrading [dSFSN23]. upstream [vdKYYWR23]. urban [vWBT20]. USA [JLX21, MB20, WBM22]. Use [Boz22, RP23, WS23]. user [Ano23b, LBCG^{+}23]. using [MG22, WB22].

v [OF22]. vaccine [YLMYY22]. valuation [Fli22, WHH24]. Value [Yeh21, YXX21, MGM^{+}21, Qia23, ST21, Wil22]. Value-Laden [YXX21]. values [Bre22, FE20]. varieties [WHL^{+}21]. variety [Ano23b, LBCG^{+}23]. various [RAM22]. vehicles [SC21b]. Venture [YKW21]. versus [BPS20, FS20, FvL22, ZXW^{+}20]. vested [CK21]. view [CCT21a, CCT21b]. views [Phi22]. virtuous [VJMHdLNA20]. vouchers [XG23]. vulnerable [SGW21].

Wales [Hen23]. Walking [SBH^{+}22]. waste [AYDQ^{+}21]. ways [JTWW21]. welfare [EZI20]. Were [AD20]. Which [ZXW^{+}20]. whom [WB22]. Wide [SLSC^{+}21, ABC^{+}21]. wide-scale [ABC^{+}21]. wind [WJZ20]. winners [DSW23]. within [NN22, SC22]. without [ASFB^{+}23]. work [Hen23, JS23, PH23, SPB20, SJ24]. Working [wY22, Ano22]. world [Rzy22, VSGMF21]. worlds [Clo21].

years [BSTG21]. young [SFP23, YCL22].

Zika [YLMYY22].

References

Akerlof:2020:NMC

[AAW^{+}20] Karen Akerlof, Alessandro Allegra, Thomas Webler, Erin Heath, Emily T. Cloyd, Carla-Leanne Washbourne, and Chris Tyler. New methods in creating transdisciplinary science policy research agendas: The case of legislative science advice. Science and Public Policy, 47 (4):536–547, August 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/4/536/5893017.

Armeni:2021:TWS

[ABC^{+}21] Kristijan Armeni, Loek Brinkman, Rickard Carlsson, Anita Eerland, Rianne Fijten, Robin Fondberg, Vera E. Heininga, Stephan Heunis, Wei Qi Koh, Maurits Masselink, Niall Moran, Andrew Ó Baoill, Alexandra Sarafoglou, Antonio Schettino, Hardy Schwamm, Zsuzsika Sjoerds, Marta Teperek, Olmo R. van den Akker, Anna van’t Veer, and Raul Zurita-Milla. Towards wide-scale adoption of open science practices: The role of open science communities. Science and Public Policy, 48 (5):605–611, October 2021. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/5/605/6313404.

AbdulRafiu:2023:PSG

[Abd23] Abbas AbdulRafiu. Policy seduction and governance resistance? Examining public funding agencies and academic institutions on decarbonisation research. Science and Public Policy, 50(1):87–101, February 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/50/1/87/6691202.

Agasisti:2021:CCA

[ABM21] Tommaso Agasisti and Jasmina Berbegal-Mirabent. Cross-country analysis of higher education institutions’ efficiency: The role of strategic positioning. Science and Public Policy, 48 (1):66–79, February 2021. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/48/1/66/6032846.

Akbaritabar:2021:INR

[ABS21] Aliakbar Akbaritabar, Giangiacomo Bravo, and Flaminio Squazzoni. The impact of a national research assessment on the publications of sociologists in Italy. Science and Public Policy, 48(5):662– 678, October 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/48/5/662/6314580.

Abramo:2020:WIP

[AD20] Giovanni Abramo and Ciriaco Andrea D’Angelo. Were the Italian policy reforms to contrast favoritism and foster effectiveness in faculty recruitment successful? Science and Public Policy, 47(5):604– 615, October 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/5/604/6026549.

Archibugi:2020:IIE

[AFF20] Daniele Archibugi, Andrea Filippetti, and Marion Frenz. Investment in innovation for European recovery: a public policy priority. Science and Public Policy, 47(1):92– 102, February 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/1/92/5618662.

Adelle:2021:CPK

[AGKL21] Camilla Adelle, Tristan Görgens, Florian Kroll, and Bruno Losch. Coproduction of knowledge in transdisciplinary communities of practice: Experiences from food governance in South Africa. Science and Public Policy, 48(2): 145–153, April 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/2/145/6025443.

Arman:2022:SID

[AIIOL22] Husam Arman, Simona Iammarino, J. Eduardo Ibarra-Olivo, and Neil Lee. Systems of innovation, diversification, and the R&D trap: a case study of Kuwait. Science and Public Policy, 49(2):179– 190, April 2022. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/49/2/179/6433072.

Adams:2022:RUC

[AJG22] Jonathan Adams, Jo Johnson, and Jonathan Grant. The rise of UK–China research collaboration: Trends, opportunities and challenges. Science and Public Policy, 49(1):132–147, February 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/49/1/132/6432140.

Alam:2022:CCI

[AKC22] Md. Razib Alam, Bonwoo Koo, and Brian Paul Cozzarin. Canada’s changing innovation landscape. Science and Public Policy, 49 (1):28–41, February 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/49/1/28/6368429.

Ahola-Launonen:2022:DEB

[ALK22] Johanna Ahola-Launonen and Sofi Kurki. Dynamics of expectations in the bioeconomy — hopes, disillusionments, and conflicting futures. Science and Public Policy, 49(6):819– 829, December 2022. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/6/819/6613523.

Acebo:2024:MIP

[AMD24] Enrique Acebo and José-Ángel Miguel-Dávila. Multilevel innovation policy mix: impact of regional, national, and European R&D grants. Science and Public Policy, 51(2):218– 235, April 2024. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/51/2/218/7329737.

Acebo:2021:IUI

[AMDN21] Enrique Acebo, José-Ángel Miguel-Dávila, and Mariano Nieto. The impact of university–industry relationships on firms’ performance: a meta-regression analysis. Science and Public Policy, 48(2):276– 293, April 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/2/276/6215148.

Anonymous:2022:CWC

[Ano22] Anonymous. Correction to: Working with curiosity: Knowledge transfer practitioners’ ambivalence at CERN. Science and Public Policy, 49(5):818, October 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/49/5/818/6726144.

See [wY22].

Anonymous:2023:CBC

[Ano23a] Anonymous. Correction to: Bridging climate change science and policy through TMNs in Turkey: CoM as a boundary object. Science and Public Policy, 50(1): 103, February 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/50/1/103/6696135.

See [EET22].

Anonymous:2023:CCC

[Ano23b] Anonymous. Correction to: Co-creating end-user roles. understanding the new variety of user involvement in public sector innovation. Science and Public Policy, 50(6): 1110, December 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/50/6/1110/7007800.

See [LBCG^{+}23].

Anonymous:2023:CEH

[Ano23c] Anonymous. Correction to:

Embracing heterogeneity: Why plural understandings strengthen interdisciplinarity and transdisciplinarity. Science and Public Policy, 50(1):102, February 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/1/102/6697109.

See [VBFLP22].

Anonymous:2023:CMS

[Ano23d] Anonymous. Correction to:

Making sense of knowledgebrokering organisations: boundary organisations or policy entrepreneurs? Science and Public Policy, 50 (6):1112, December 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/6/1112/7255893.

See [MCDD23].

Anonymous:2023:COS

[Ano23e] Anonymous. Correction to:

Open science-related policies in Europe. Science and Public Policy, 50(6): 1111, December 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/6/1111/7109885.

See [MA23].

Atta-Owusu:2022:WMA

[AOF22] Kwadwo Atta-Owusu and Rune Dahl Fitjar. What motivates academics for external engagement? Exploring the effects of motivational drivers and organizational fairness. Science and Public Policy, 49(2): 201–218, April 2022. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/2/201/6430243.

Adelle:2020:MST

[APGL20] Camilla Adelle, Laura Pereira, Tristan Görgens, and Bruno Losch. Making sense together: The role of scientists in the coproduction of knowledge for policy making. Science and Public Policy, 47(1): 56–66, February 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/1/56/5622756.

Ayoubi:2021:DIP

[APV21] Charles Ayoubi, Michele Pezzoni, and Fabiana Visentin. Does it pay to do novel science? The selectivity patterns in science funding. Science and Public Policy, 48(5):635– 648, October 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/5/635/6296161.

Alaja:2020:ENI

[AS20] Antti Alaja and Ville-Pekka Sorsa. The evolution of the national innovation system as programmatic policy idea in Finland. Science and Public Policy, 47(6):834–843, December 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/6/834/5942750.

AssisFeitosa:2023:DIR

[ASFB^{+}23] Paulo Henrique Assis Feitosa, Sergio Salles-Filho, Adriana Bin, Yohanna Juk, and Fernando Antonio Basile Colugnati. Does international R&D cooperation under institutional agreements have a greater impact than those without agreements? Science and Public Policy, 50(5):831– 841, October 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/5/831/7181384.

Abma-Schouten:2023:ERP

[ASGRM23] Rebecca Abma-Schouten, Joey Gijbels, Wendy Reijmerink, and Ingeborg Meijer. Evaluation of research proposals by peer review panels: broader panels for broader assessments? Science and Public Policy, 50 (4):619–632, August 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/50/4/619/7115831.

Alves:2021:MOP

[AVZ21] André C. Alves, Nicholas S. Vonortas, and Paulo A. Zawislak. Mission-oriented policy for innovation and the fuzzy boundary of market creation: the Brazilian shipbuilding case. Science and Public Policy, 48(1): 80–92, February 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/48/1/80/6032886.

Adobor:2023:PAI

[AY23] Henry Adobor and Robert Yawson. The promise of artificial intelligence in combating public corruption in the emerging economies: a conceptual framework. Science and Public Policy, 50(3):355–370, June 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/50/3/355/6889870.

Akon-Yamga:2021:TIP

[AYDQ^{+}21] Gordon Akon-Yamga, Chux U. Daniels, Wilhemina Quaye, Blanche M. Ting, and Adelaide A. Asante. Transformative innovation policy approach to e-waste management in Ghana: Perspectives of actors on transformative changes. Science and Public Policy, 48(3): 387–397, June 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/3/387/6141578.

Bartroli:2023:NSR

[Bar23] Montserrat Alom Bartrolí.

North–South research funding dynamics of collaborative projects: researchers’ appropriation strategies of agencies’ project frameworks. Science and Public Policy, 50(5):935–946, October 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/5/935/7202184.

Baveye:2023:PHB

[Bav23] Philippe C. Baveye. Perspective on honest brokers in science: Broadening conflicts of interest and role academies could play. Science and Public Policy, 50(1):129–132, February 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/50/1/129/6718168.

Barazzetti:2021:MPM

[BBAK21] Gaia Barazzetti, Nolwenn Bühler, Marc Audétat, and Alain Kaufmann. Making personalized medicine ethical: a critical examination of the new promises of ‘personalized health’ in Switzerland. Science and Public Policy, 48(6):818– 828, December 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/48/6/818/6340522.

Bauer:2023:DIO

[BCFU23] Anja Bauer, Leo Capari, Daniela Fuchs, and Titus Udrea. Diversification, integration, and opening: developments in modelling for policy. Science and Public Policy, 50(6):977– 987, December 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/50/6/977/7233825.

Benner:2022:IPS

[Ben22] Maximilian Benner. An institutionalist perspective on smart specialization: Towards a political economy of regional innovation policy. Science and Public Policy, 49(6):878– 889, December 2022. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/6/878/6619344.

Binder:2020:IRS

[BFH^{+}20] Claudia R. Binder, Livia Fritz, Ralph Hansmann, Andreas Balthasaar, and Zilla Roose. Increasing the relevance of science for practice and practice for science: Quantitative empirical insights. Science and Public Policy, 47(6):772– 787, December 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/6/772/6048945.

Bottcher:2022:RAR

[BGW22] Lucas Böttcher, Hans Gersbach, and Didier Wernli. Restoring the antibiotic R&D market to combat the resistance crisis. Science and Public Policy, 49(1): 127–131, February 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/1/127/6396037.

See erratum [BWG22].

Baek:2022:RPA

[BL22] Seung Chul Baek and Duk Hee Lee. Role of professional autonomy and project commitment in the Korean government-funded R&D projects. Science and Public Policy, 49(1): 1–17, February 2022. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/49/1/1/6366389.

Blair:2021:EST

[Bla21a] Peter D. Blair. Effective science and technology assessment advice for congress: comparing options. Science and Public Policy, 48(2):164–176, April 2021. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/48/2/164/6067289.

See erratum [Bla21b].

Blair:2021:EES

[Bla21b] Peter D. Blair. Erratum to: Effective science and technology assessment advice for Congress: comparing options. Science and Public Policy, 48(2): 294, April 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/48/2/294/6173808.

See [Bla21a].

Brett:2023:GCG

[BMA23] Nancy Brett, Thomas Magnusson, and Hans Andersson. From global climate goals to local practice — mission-oriented policy enactment in three Swedish regions. Science and Public Policy, 50(4): 603–618, August 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/4/603/7111226.

Bonaccorsi:2021:CRI

[BMCF21] Andrea Bonaccorsi, Nicola Melluso, Filippo Chiarello, and Gualtiero Fantoni. The credibility of research impact statements: a new analysis of REF with semantic hypergraphs. Science and Public Policy, 48 (2):212–225, April 2021. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/2/212/6132801.

Borsi:2021:BSA

[Bor21] Balázs Borsi. The balanced state of applicationoriented public research and technology organisations. Science and Public Policy, 48(5):612–629, October 2021. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/5/612/6320915.

Bozeman:2022:USP

[Boz22] Barry Bozeman. Use of science in public policy: Lessons from the COVID-19 pandemic efforts to ‘follow the science’. Science and Public Policy, 49 (5):806–817, October 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/49/5/806/6603992.

Boardman:2020:SPI

[BPS20] Craig Boardman, Branco Ponomariov, and H. L. Shen. Science policy implementation: strategic versus technical human resource management. Science and Public Policy, 47(6):844– 854, December 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/47/6/844/6043727.

Brenninkmeijer:2022:ASA

[Bre22] Jonna Brenninkmeijer. Achieving societal and academic impacts of research: a comparison of networks, values, and strategies. Science and Public Policy, 49 (5):728–738, October 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/49/5/728/6585532.

Bugge:2021:EPM

[BS21] Markus M. Bugge and Fazilat Siddiq. Empowering professionalism in missionoriented innovation. Science and Public Policy, 48(3):423–437, June 2021. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/3/423/6252980.

Bianchi:2022:CTU

[BS22a] Federico Bianchi and Flaminio Squazzoni. Can transparency undermine peer review? A simulation model of scientist behavior under open peer review. Science and Public Policy, 49 (5):791–800, October 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/5/791/6602348.

Borras:2022:DTR

[BS22b] Susana Borrás and Sylvia Schwaag Serger. The design of transformative research and innovation policy instruments for grand challenges: The policy-nesting perspective. Science and Public Policy, 49(5):659–672, October 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/5/659/6554530.

Bastos:2021:FYU

[BSTG21] Elisa Cordeiro Bastos, Aline Rossales Sengik, and Jorge Tello-Gamarra. Fifty years of university– industry collaboration: a global bibliometrics overview. Science and Public Policy, 48(2):177–199, April 2021. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/48/2/177/6163153.

Banozic-Tang:2022:PRP

[BTT22] Adriana Banozic-Tang and Araz Taeihagh. Perspective on research–policy interface as a partnership: The study of best practices in CREATE. Science and Public Policy, 49(5):801– 805, October 2022. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/49/5/801/6604793.

Bottcher:2022:ERA

[BWG22] Lucas Böttcher, Didier Wernli, and Hans Gersbach. Erratum to: Restoring the antibiotic R&D market to combat the resistance crisis. Science and Public Policy, 49(2): 267, April 2022. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/

article/49/2/267/6461727.

See [BGW22].

Braams:2022:UWC

[BWMH22] Rik B. Braams, Joeri H.

Wesseling, Albert J. Meijer, and Marko P. Hekkert. Understanding why civil servants are reluctant to carry out transition tasks. Science and Public Policy, 49(6):905–914, December 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/6/905/6644959.

Bo:2021:ESP

[BZG21] Qingwen Bo, Bo Zou, and Feng Guo. Environmental scanning perspective: The moderating effects of organizational risk and government subsidy in innovation search process. Science and Public Policy, 48 (4):531–540, August 2021. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/4/531/6289862.

Cavallaro:2024:UAS

[Cav24] Marco Cavallaro. Universities of applied sciences’ EU research project participation through the lens of differentiation. Science and Public Policy, 51(1):1– 14, February 2024. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/51/1/1/7246612.

Caliari:2023:BAF

[CBUS23] Thiago Caliari, Mayara Bovo, Lígia Maria Soto Urbina, and Rodrigo Arnaldo Scarpel. Brazilian Air Force acquisition policies: observing absorptive capacity and contingent factors in aeronautical beneficiary companies. Science and Public Policy, 50 (5):917–934, October 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/50/5/917/7202183.

Cao:2020:RSE

[CBWJ20] Cong Cao, Jeroen Baas, Caroline S. Wagner, and Koen Jonkers. Returning scientists and the emergence of China’s science system. Science and Public Policy, 47(2):172–183, April 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/47/2/172/5658550.

Coco:2024:BCF

[CCF24] Nunzia Coco, Cinzia Colapinto, and Vladi Finotto. Bridging conflicting frames in policies for digital transformation. Science and Public Policy, 51(2):285– 296, April 2024. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/51/2/285/7412280.

Chang:2021:EHD

[CCT21a] Yuan-Chieh Chang, Po-Hsuan Chen, and Min-Jun Teng. Erratum to: How do institutional changes facilitate university-centric networks in Taiwan? The Triple-Helix model of innovation view. Science and Public Policy, 48(4):602– 603, August 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/4/602/6318635.

See [CCT21b].

Chang:2021:HDI

[CCT21b] Yuan-Chieh Chang, Po-Hsuan Chen, and Min-Jun Teng. How do institutional changes facilitate university-centric networks in Taiwan? The Triple Helix model of innovation view. Science and Public Policy, 48(3): 309–324, June 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/

                                        article/48/3/309/6291354.

See erratum [CCT21a].

Cho:2020:EUP

[CH20a] Seongkyung Cho and Christopher S. Hayter. Erratum to: Under pressure: a systematic review of stress and its impact among graduate students. Science and Public Policy, 47(6): 893, December 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/47/6/893/6179537.

See [CH20b].

Cho:2020:UPS

[CH20b] Seongkyung Cho and Christopher S. Hayter. Under pressure: a systematic review of stress and its impact among graduate students. Science and Public Policy, 47(6): 758–771, December 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/47/6/758/6025444.

See erratum [CH20a].

Chang:2022:TSN

[Cha22] Shu-Hao Chang. Technological structure network analysis to explore the hotspots of academic patents in international technology transfer. Science and Public Policy, 49 (1):98–114, February 2022.

CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/1/98/6366461.

Chavez:2024:UCT

[CHH^{+}24] Cecilia Garcia Chavez, David Howoldt, Patrick Hoyer, Maria Karaulova, Henning Kroll, and Torben Schubert. Understanding career transitions of applied researchers to universities: evidence from Germany. Science and Public Policy, 51(1):89– 107, February 2024. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/51/1/89/7282747.

Choi:2021:OVI

[CK21] Yong In Choi and Byoung-Joo Kim. Overcoming vested interests against innovation through political entrepreneurship: a comparative study of Korean mobility cases. Science and Public Policy, 48(3): 412–422, June 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/3/412/6294473.

Cocos:2020:WWK

[CL20] Marta Cocos and Benedetto Lepori. What we know about research policy mix.

Science and Public Policy, 47(2):235–245, April 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/47/2/235/5714941.

Cohen:2024:ESC

[CLiM^{+}24] Joshua B. Cohen, Anne M. C. Loeber, ilse Marschalek, Michael J. Bernstein, Vincent Blok, Raúl Tabarés, Robert Gianni, and Erich Griessler. From experimentation to structural change: fostering institutional entrepreneurship for public engagement in research and innovation. Science and Public Policy, 51(2):324– 336, April 2024. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/51/2/324/7423039.

Clormann:2021:SBW

[Clo21] Michael Clormann. Switching between worlds apart: Negotiating European space sector cultures through innovation. Science and Public Policy, 48(4):521– 530, August 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/48/4/521/6278461.

Chowdhury:2022:SDP

[CLvH22] Farhat Chowdhury, Albert N. Link, and Martijn van Hasselt. The spatial distribution of public support for AI research. Science and Public Policy, 49 (4):573–579, August 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/4/573/6540150.

Cevallos:2020:NPC

[CM20] Rodrigo A. Cevallos and Carlos Merino Moreno. National policy councils for science, technology, and innovation: a scheme for structural definition and implementation. Science and Public Policy, 47(5): 705–718, October 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/5/705/6055889.

Coombs:2021:TER

[CM21] Sarah K. Coombs and Ingeborg Meijer. Towards evaluating the research impact made by universities of applied sciences. Science and Public Policy, 48(2): 226–234, April 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/2/226/6140182.

Christensen:2020:WNPa

[CNA^{+}20a] Malene Vinther Christensen, Mika Nieminen, Marlene Altenhofer, Elise Tancoigne, Niels Mejlgaard Erich Griessler, and Adolf Filacek. What’s in a name? Perceptions and promotion of responsible research and innovation practices across Europe. Science and Public Policy, 47(3):360–370, June 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/3/360/5834671.

See author name correction [CNA^{+}20b] and [CNA^{+}20a].

Christensen:2020:CWN

[CNA^{+}20b] Malene Vinther Christensen, Mika Nieminen, Marlene Altenhofer, Elise Tancoigne, Niels Mejlgaard, Erich Griessler, and Adolf Filacek. Correction: What’s in a name? Perceptions and promotion of responsible research and innovation practices across Europe. Science and Public Policy, 47(6): 892, December 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/6/892/5918209.

See [CNA^{+}20a, CNA^{+}20c].

Christensen:2020:WNPb

[CNA^{+}20c] Malene Vinther Christensen, Mika Nieminen, Marlene Altenhofer, Elise Tangcoigne, Niels Mejlgaard, Erich Griessler, and Adolf Filacek. What’s in a name? Perceptions and promotion of responsible research and innovation practices across Europe. Science and Public Policy, 47(6):892, December 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/6/892/5918209.

See [CNA^{+}20a, CNA^{+}20b].

Cohen:2022:IPE

[Coh22] Joshua B. Cohen. Institutionalizing public engagement in research and innovation: Toward the construction of institutional entrepreneurial collectives. Science and Public Policy, 49(5):673–685, October 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/5/673/6561928.

Choi:2023:GGI

[CP23] Hyeri Choi and Min Jae Park. To govern or be governed: an integrated framework for AI governance in the public sector. Science and Public Policy, 50(6): 1059–1072, December 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/50/6/1059/7239087.

Cacace:2024:CCC

[CPT^{+}24] Marina Cacace, Francesca Pugliese, Charikleia Tzanakou, Jörg Müller, Alain Denis, and Maria Sangiuliano. Certifying complexity? The case of a European gender equality certification scheme for research-performing organizations. Science and Public Policy, 51(2):261– 273, April 2024. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/51/2/261/7343993.

Cheng:2020:DPI

[CZL^{+}20] Hua Cheng, Zhiying Zhang, Zhongju Liao, Yong Wei, and Joseph Martial Nkongo Mvondo. Different policy instruments and the threshold effects on collaboration efficiency in China. Science and Public Policy, 47(3): 348–359, June 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/47/3/348/5807894.

Conle:2021:TTM

[CZtB21] Marcus Conlé, Wei Zhao, and Tobias ten Brink. Technology transfer models for knowledge-based regional development: New R&D institutes in Guangdong, China. Science and Public Policy, 48(1):132– 144, February 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/1/132/6032885.

DePeuter:2023:FRI

[DC23] Steven De Peuter and Stijn Conix. Fostering a research integrity culture: Actionable advice for institutions. Science and Public Policy, 50(1):133– 145, February 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/1/133/6749026.

Doloreux:2022:ISI

[DF22] David Doloreux and Anthony Frigon. The Innovation Superclusters Initiative in Canada: a new policy strategy? Science and Public Policy, 49(1):148– 158, February 2022. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/1/148/6432302.

deJong:2022:HUI

[dJB22] Stefan P. L. de Jong and Corina Balaban. How universities influence societal impact practices: Academics’ sense-making of organizational impact strategies. Science and Public Policy, 49(4):609–620, August 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/4/609/6552261.

deJong:2022:PIE

[dJBN22] Stefan P. L. de Jong, Corina Balaban, and Maria Nedeva. From ‘productive interactions’ to ‘enabling conditions’: The role of organizations in generating societal impact of academic research. Science and Public Policy, 49(4): 643–645, August 2022. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/4/643/6552260.

delaGuardia:2021:FOP

[dlGGM21] Fernando Hoces de la Guardia, Sean Grant, and Edward Miguel. A framework for open policy analysis. Science and Public Policy, 48(2):154–163, April 2021. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/2/154/6018524.

delaTorre:2021:KTP

[dlTGRS21] Eva M. de la Torre, Maryam Ghorbankhani, Federica Rossi, and Marti Sagarra. Knowledge transfer profiles of public research organisations: the role of fields of knowledge specialisation. Science and Public Policy, 48(6):860– 876, December 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/6/860/6355300.

Silva:2023:BMI

[dMSVF23] Diego R. de Moraes Silva, Nicholas S. Vonortas, and André T. Furtado. Barriers as moderators in the innovation process. Science and Public Policy, 50(4): 633–654, August 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/4/633/7140305.

Donadelli:2020:WED

[Don20] Flavia Donadelli. When evidence does not matter: The barriers to learning from science in two cases of environmental policy change in Brazil. Science and Public Policy, 47(3):313–321, June 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/47/3/313/5734686.

Dusdal:2021:BMC

[DP21] Jennifer Dusdal and Justin J. W. Powell. Benefits, motivations, and challenges of international collaborative research: a sociology of science case study. Science and Public Policy, 48(2): 235–245, April 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/48/2/235/6135106.

deRassenfosse:2020:SKF

[dRS20] Gaétan de Rassenfosse and Florian Seliger. Sources of knowledge flow between developed and developing nations. Science and Public Policy, 47(1):16– 30, February 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/47/1/16/5580327.

Dutfield:2023:OIS

[DS23] Graham Dutfield and Katerina Sideri. Openness, innovation, and science policy in the age of datadriven medicine. Science and Public Policy, 50(6): 947–949, December 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/6/947/5818954.

daSilva:2023:ITM

[dSFSN23] Renan G. L. da Silva, Bruno B. Fischer, Paola R. Schaeffer, and Hillegonda M. D. Novaes. The industry of therapeutic monoclonal antibodies in Brazil: Public policies as instruments of technology upgrading. Science and Public Policy, 50 (1):42–58, February 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/1/42/6686552.

Teixeira:2020:ODI

[dSTRC20] André Luiz da Silva Teixeira, Márcia Siqueira Rapini, and Thiago Caliari. Organizational determinants and idiosyncrasies of firms’ absorptive capacity in a developing country. Science and Public Policy, 47(3): 384–395, June 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/3/384/5843890.

Denney:2023:DWP

[DSW23] Steven Denney, Travis Southin, and David A. Wolfe. Do winners pick government? How scaleup experience shapes entrepreneurs’ assessments of innovation policy mixes.

Science and Public Policy, 50(5):858–870, October 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/50/5/858/7197240.

Duiveman:2020:IER

[Dui20] Robert (R. M. ) Duiveman. Institutionally embedded research; how academic and governmental practices enable and constrain collaborative research. Science and Public Policy, 47(6): 855–864, December 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/47/6/855/6100898.

Donati:2023:LAS

[DWK23] Letizia Donati and Caroline Wigren-Kristoferson. A legitimacy approach to social innovation initiatives at universities. Science and Public Policy, 50 (2):194–205, April 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

                                        https://academic.oup.
                                        com/spp/article/50/2/
                                        194/6783007.

Erbil:2022:BCC

[EET22] Aslı Öğüt Erbil, Mehmet Eroğlu, and Ebru Gönül Türk. Bridging climate change science and policy through TMNs in Turkey: CoM as a boundary-object. Science and Public Policy, 49(6):830–842, December 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/6/830/6640332.

See correction [Ano23a].

Enger:2020:OCP

[EG20] Simen G. Enger and Magnus Gulbrandsen. Orchestrating collaborative projects: Inside ICT networks in Horizon 2020. Science and Public Policy, 47(3):396–409, June 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/3/396/5828389.

Ehsan:2021:BIS

[Ehs21] Firoz Ehsan. Boosting innovation in small- and medium-sized enterprises through tax incentives: lessons from the UK. Science and Public Policy, 48 (5):712–726, October 2021. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/5/712/6328884.

Echeverria-King:2023:SCB

[EKFR^{+}23] Luisa F. Echeverría-King, Aura Fossati, Nussaïbah B.

Raja, Kleinsy Bonilla, Bernardo Urbani, Radenka Krsmanović Whiffen, and Tereza Vizinová. Scientific collaborations between Latin America and Europe: an approach from science diplomacy towards international engagement. Science and Public Policy, 50(4): 794–806, August 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/4/794/7190065.

Edler:2024:SIT

[EOS24] Jakob Edler, Katrin Ostertag, and Johanna Schuler. Social innovation, transformation, and public policy: towards a conceptualization and critical appraisal. Science and Public Policy, 51(1):80–88, February 2024. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/51/1/80/7281956.

Edquist:2020:FPI

[EZI20] Charles Edquist and Jon Mikel Zabala-Iturriagagoitia. Functional procurement for innovation, welfare, and the environment. Science and Public Policy, 47(5):595– 603, October 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/5/595/6026548.

Ferreira:2024:PPP

[FD24] António Ferreira and João Quesado Delgado. The precarity paradox: the precaritydriven inefficiencies of research at a public university. Science and Public Policy, 51(2):297–308, April 2024. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/51/2/297/7420770.

Feldbacher-Escamilla:2020:KVR

[FE20] Christian J. Feldbacher-Escamilla. Knowledge and values: a re-entanglement in epistemic regimes. Science and Public Policy, 47 (1):67–77, February 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/1/67/5613037.

Fecher:2023:BIB

[FKF^{+}23] Benedikt Fecher, Freia Kuper, Birte Fähnrich, Hannah Schmid-Petri, Thomas Schildhauer, Peter Weingart, and Holger Wormer. Balancing interests between freedom and censorship: Organizational strategies for quality assurance in science communication. Science and Public Policy, 50(1):1–14, February 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/50/1/1/6658234.

Fecher:2021:MRI

[FKS^{+}21] Benedikt Fecher, Rebecca Kahn, Nataliia Sokolovska, Teresa Völker, and Philip Nebe. Making a research infrastructure: Conditions and strategies to transform a service into an infrastructure. Science and Public Policy, 48(4):499– 507, August 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/48/4/499/6250984.

Fiske:2023:RUC

[FLBM23] Amelia Fiske, Johannes Lange, Alena Buyx, and Stuart McLennan. Responding to uncertainty in the COVID-19 pandemic: perspectives from Bavaria, Germany. Science and Public Policy, 50(5):881– 892, October 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/50/5/881/7197919.

Flink:2022:TPS

[Fli22] Tim Flink. Taking the pulse of science diplomacy and developing practices of valuation. Science and Public Policy, 49(2):191– 200, April 2022. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/2/191/6430235.

Fabiano:2020:CPK

[FMF20] Gianluca Fabiano, Andrea Marcellusi, and Giampiero Favato. Channels and processes of knowledge transfer: How does knowledge move between university and industry? Science and Public Policy, 47(2): 256–270, April 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/2/256/5721499.

Falk:2020:ECV

[FS20] Martin Thomas Falk and Roger Svensson. Evaluation criteria versus firm characteristics as determinants of public R&D funding. Science and Public Policy, 47(4):525–535, August 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/4/525/5900380.

Ferretti:2022:PMM

[FvL22] Federico Ferretti and Harro van Lente. The promise of the Maker Movement: policy expectations versus community criticisms. Science and Public Policy, 49 (1):18–27, February 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/1/18/6414640.

Fan:2022:TPC

[FZW22] Fei Fan, Xuerong Zhang, and Xueli Wang. Are there political cycles hidden inside collaborative innovation efficiency? An empirical study based on Chinese cities. Science and Public Policy, 49(3):532– 551, June 2022. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/3/532/6545201.

Garcia-Carbonell:2021:AHC

[GCGAMASG21] Natalia García-Carbonell, Félix Guerrero-Alba, Fernando Martín-Alcázar, and Gonzalo Sánchez-Gardey. Academic human capital in universities: definition and proposal of a measurement scale. Science and Public Policy, 48(6):877– 888, December 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/6/877/6355307.

Gunn:2023:CDC

[GFF^{+}23] Callum J. Gunn, Sevgi E.

Fruytier, Teresa Finlay, Lidewij Eva Vat, Teun Zuiderent-Jerak, and Tjerk Jan Schuitmaker-Warnaar. Codesign and its consequences: developing a shared patient engagement framework in the IMI-PARADIGM project. Science and Public Policy, 50(6):1018–1028, December 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/6/1018/7235647.

Ghosh:2020:MPM

[Gho20] Shubha Ghosh. Myriad post-Myriad. Science and Public Policy, 47(5):638– 646, October 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/5/638/5766083.

Ghosh:2021:TOA

[GKR^{+}21] Bipashyee Ghosh, Paula Kivimaa, Matias Ramirez, Johan Schot, and Jonas Torrens. Transformative outcomes: assessing and reorienting experimentation with transformative innovation policy. Science and Public Policy, 48(5):739– 756, October 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/48/5/739/6332811.

Graddy-Reed:2021:DIC

[GRFBL21] Alexandra Graddy-Reed, Maryann Feldman, Janet Bercovitz, and W. Scott Langford. The distribution of indirect cost recovery in academic research. Science and Public Policy, 48(3): 364–386, June 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/48/3/364/6199575.

Graddy-Reed:2023:PDA

[GRL23] Alexandra Graddy-Reed and Lauren Lanahan. Prioritizing diversity? The allocation of US federal R&D funding. Science and Public Policy, 50(1):104– 119, February 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/50/1/104/6705406.

Ghazinoory:2024:ESL

[GRS24] Sepehr Ghazinoory, Alireza Ranjbar, and Tahereh Sonia Saheb. Evolution of the STIP literature: discovering the growing role of innovation governance concepts. Science and Public Policy, 51(1):162– 177, February 2024. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/51/1/162/7319532.

Garcia-Sanchez:2022:CIC

[GSR22] Antonio García-Sánchez and Ruth Rama. Cooperative innovation and crises: Foreign subsidiaries, state-owned enterprises, and domestic private firms. Science and Public Policy, 49(6):915–927, December 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/6/915/6645283.

Grimpe:2020:CIR

[GST^{+}20] Barbara Grimpe, Bernd Carsten Stahl, Carolyn Ten Holter, Philip Inglesant, Grace Eden, Menisha Patel, and Marina Jirotka. From collaborative to institutional reflexivity: Calibrating responsibility in the funding process. Science and Public Policy, 47(5):720– 732, October 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/5/720/5924402.

Gomez-Valenzuela:2020:SPD

[GV20] Víctor Gómez-Valenzuela.

STI policies in the Dominican Republic: the influence of economic rationales from a contextdevelopment perspective. Science and Public Policy, 47(3):371–383, June 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/47/3/371/5830821.

Goser:2023:RES

[GWS23] Maya Göser, Stefan Wimmer, and Johannes Sauer. Research excellence and scientific advisory boards. Science and Public Policy, 50(6):1044–1058, December 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/50/6/1044/7236569.

Guo:2023:HPR

[GZG^{+}23] Jianfeng Guo, Xuemei Zhang, Fu Gu, Jiannan Zhu, Chao Deng, Xinze Zhao, and Xiaohan Yang. Herding in policy responses to coronavirus disease 2019. Science and Public Policy, 50(5):893–904, October 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/50/5/893/7199924.

Hesjedal:2023:MST

[HÅ23] Maria Bårdsen Hesjedal and Heidrun Åm. Making sense of transdisciplinarity: Interpreting science policy in a biotechnology centre. Science and Public Policy, 50(2):219–229, April 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

https://academic.oup.
com/spp/article/50/2/
219/6809965.

Hawkins:2020:PNI

[Haw20] Naomi Hawkins. Patents and non-invasive prenatal testing: Is there cause for concern? Science and Public Policy, 47(5):655– 667, October 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/5/655/5828386.

Hemphill:2023:TEI

[Hem23] Thomas A. Hemphill. Technology entrepreneurship and innovation hubs: Perspectives on the universal regulatory sandbox. Science and Public Policy, 50 (2):350–353, April 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

https://academic.oup.
com/spp/article/50/2/
350/6889949.

Henderson:2023:BWR

[Hen23] Dylan Henderson. Boundary work in the regional innovation policy mix: SME digital technology diffusion policies in Wales. Science and Public Policy, 50(3): 548–558, June 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/50/3/548/7077427.

Herberholz:2023:NFC

[Her23] Lars Herberholz. A new facet of cumulative advantage in higher education finance. Science and Public Policy, 50(5):819– 830, October 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/50/5/819/7095647.

Hennen:2021:CMR

[HHL21] Leonhard Hennen, Julia Hahn, and Miltos Ladikas. Competing modes of responsibility in research organisations — insights from an international comparison. Science and Public Policy, 48(1):54–65, February 2021. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/48/1/54/6044343.

Huggins:2023:COI

[HJMX23] Robert Huggins, Andrew Johnston, Max Munday, and Chen Xu. Competition, open innovation, and growth challenges in the semiconductor industry: the case of Europe’s clusters. Science and Public Policy, 50(3): 531–547, June 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/3/531/7072258.

Ku:2022:DII

[hKZ22] Sharon Tsai hsuan Ku and Stephen Zehr. Disciplining interdisciplinarity: Infrastructure, identity, and interdisciplinary practice in nanoELSI research. Science and Public Policy, 49 (5):765–780, October 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/5/765/6596171.

Hayter:2020:GME

[HL20] Christopher S. Hayter and Albert N. Link. Governance mechanisms enabling inter-organizational adaptation: Lessons from grand challenge R&D programs. Science and Public Policy, 47(2):271–282, April 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/2/271/5721458.

Hu:2021:TKL

[HL21] Albert G. Z. Hu and Vu Thinh Ly. Tacit knowledge, localization push, and diffusion of science: Evidence from Singapore. Science and Public Policy, 48 (1):16–26, February 2021. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/1/16/5989818.

Hottenrott:2022:WBM

[HL22] Hanna Hottenrott and Cornelia Lawson. What is behind multiple institutional affiliations in academia? Science and Public Policy, 49(3):382–402, June 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/3/382/6461995.

Hastings:2023:GRI

[HLL^{+}23] Rosie Hastings, Krishma Labib, Iris Lechner, Lex Bouter, Guy Widdershoven, and Natalie Evans. Guidance on research integrity provided by pan-European discipline-specific learned societies: a scoping review. Science and Public Policy, 50(2):318– 335, April 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

https://academic.oup.
com/spp/article/50/2/
318/6881910.

Huang:2023:DPC

[HLM^{+}23] Ying Huang, Yashan Li, Jinge Mao, Ruinan Li, and Lin Zhang. The dynamics of policy coordination: The case of China’s science and technology policy-making. Science and Public Policy, 50(2):177–193, April 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

https://academic.oup.
com/spp/article/50/2/
177/6762876.

Hansen:2020:CMR

[HLS20] Hanne Foss Hansen, Jonas Krog Lind, and Andreas Kjær Stage. Changing managerial roles in Danish universities. Science and Public Policy, 47(4):447– 457, August 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/4/447/5837959.

Petersen:2022:SUT

[hPKvR22] Il haam Petersen, Glenda Kruss, and Nicole van Rheede. Strengthening the university third mission through building community capabilities alongside university capabilities. Science and Public Policy, 49(6):890–904, December 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/49/6/890/6639597.

Harmsen:2022:SHR

[HPV^{+}22] Simone Harmsen, Carina A. C. M. Pittens, Eva Vroonland, Annemiek J. M. L. van Rensen, and Jacqueline E. W. Broerse. Supporting health researchers to realize meaningful patient involvement in research: Exploring researchers’ experiences and needs. Science and Public Policy, 49(5):751–764, October 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/49/5/751/6583406.

Hansen:2020:CAE

[HR20] Thomas Trøst Hansen and Carina Ren. Chairs of academic events: The investments and academic impact. Science and Public Policy, 47(3):322– 332, June 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/47/3/322/5732865.

Horbach:2023:DLC

[HSAR23] Serge P. J. M. Horbach, Mads P. Sørensen, Nick Allum, and Abigail-Kate Reid. Disentangling the local context — imagined communities and researchers’ sense of belonging. Science and Public Policy, 50(4):695–706, August 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/4/695/7161514.

Herberg:2021:BSS

[HSSS21] Jeremias Herberg, Seán Schmitz, Dorota Stasiak, and Gregor Schmieg. Boundary speak in sustainability studies: Computational reading of a transversal field. Science and Public Policy, 48(3):398– 411, June 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/3/398/6295737.

Hu:2021:MMI

[HSYW21] Die Hu, Maoyan She, Lingfeng Ye, and Zhiwei Wang. The more the merrier? Inventor team size, diversity, and innovation quality. Science and Public Policy, 48(4):508– 520, August 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/4/508/6263441.

Horbach:2022:PLC

[HTB22] Serge P. J. M. Horbach, Joeri K. Tijdink, and Lex M. Bouter. Partial lottery can make grant allocation more fair, more efficient, and more diverse. Science and Public Policy, 49(4):580– 582, August 2022. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/49/4/580/6542962.

Haglin:2023:IKA

[HV23] Kathryn Haglin and Arnold Vedlitz. Ideology, knowledge, and the assessment of science policy agencies. Science and Public Policy, 50 (4):707–718, August 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/50/4/707/7188131.

Hwang:2021:CAR

[HYO^{+}21a] Won-Sik Hwang, Yeongjun Yeo, Inha Oh, Chanyoung Hong, Sungmoon Jung, Heewon Yang, and Jeong-Dong Lee. CGE analysis of R&D investment policy considering trade-offs between economic growth and stability. Science and Public Policy, 48(3):295– 308, June 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/3/295/6050627.

See corrigendum [HYO^{+}21b].

Hwang:2021:CCA

[HYO^{+}21b] Won-Sik Hwang, Yeongjun Yeo, Inha Oh, Chanyoung Hong, Sungmoon Jung, Heewon Yang, and Jeong-Dong Lee. Corrigendum to: CGE analysis of R&D investment policy considering trade-offs between economic growth and stability. Science and Public Policy, 48(3):445, June 2021. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/3/445/6231617.

See [HYO^{+}21a].

Ireni-Saban:2022:CER

[ISS22] Liza Ireni-Saban and Maya Sherman. Cyborg ethics and regulation: ethical issues of human enhancement. Science and Public Policy, 49(1):42–53, February 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/1/42/6366837.

Ito:2020:MAR

[IW20] Shin Ito and Toshiya Watanabe. Multilevel analysis of research management professionals and external funding at universities: Empirical evidence from Japan. Science and Public Policy, 47(6):747– 757, December 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/6/747/6104052.

Jonsson:2022:UKC

[JGK22] Anna Jonsson, Maria Grafström, and Mikael Klintman. Unboxing knowledge in collaboration between academia and society: a story about conceptions and epistemic uncertainty. Science and Public Policy, 49(4):583– 597, August 2022. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/4/583/6543972.

Joly:2021:PFT

[JL21] Pierre-Benoit Joly and Claire Le Renard. The past futures of techno-scientific promises. Science and Public Policy, 48(6):900– 910, December 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/6/900/6357827.

Jiang:2021:EBT

[JLX21] Nan Jiang, Xing Liu, and Ming Xu. Evaluating blockchain technology and related policies in China and the USA. Science and Public Policy, 48(4):562– 575, August 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/4/562/6291644.

Jorgensen:2023:NWE

[JS23] Susanne Jørgensen and Knut H. Sørensen. Numeric work: The efforts of calculation actors to make numbers count in climate and energy policy. Science and Public Policy, 50(2):243–252, April 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

https://academic.oup.
com/spp/article/50/2/
243/6825183.

Janssen:2021:PPM

[JTWW21] Matthijs J. Janssen, Jonas Torrens, Joeri H. Wesseling, and Iris Wanzenböck. The promises and premises of mission-oriented innovation policy — a reflection and ways forward. Science and Public Policy, 48(3): 438–444, June 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/3/438/6298315.

Jiang:2022:LPT

[JWFY22] Zhenyu Jiang, Zongjun Wang, Chengxiao Feng, and Lin Yi. Local political turnover, R&D investment leap and corporate innovation performance: Evidence from China. Science and Public Policy, 49(2): 347–364, April 2022. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/2/347/6519187.

Janssen:2023:MBO

[JWT^{+}23] Matthijs J. Janssen, Joeri Wesseling, Jonas Torrens, K. Matthias Weber, Caetano Penna, and Laurens Klerkx. Missions as boundary objects for transformative change: understanding coordination across policy, research, and stakeholder communities. Science and Public Policy, 50(3):398–415, June 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/3/398/7030831.

Kim:2020:OCT

[KCL20] Kyunam Kim, Chung Choe, and Daeho Lee. Overlapping certification and technical efficiency of ICT convergence companies in South Korea. Science and Public Policy, 47 (4):514–524, August 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/4/514/5903516.

Kim:2022:REM

[KH22] Dohyoung Kim and Junseok Hwang. Is renewable energy more favorable to diversity than conventional energy sources on R&D performance? Science and Public Policy, 49 (4):646–658, August 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/4/646/6567486.

Kim:2021:PII

[Kim21] Jongheon Kim. Promoting the ICT industry for the future with fears from the past. Science and Public Policy, 48(6):889– 899, December 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/6/889/6355930.

Kishi:2020:HDP

[Kis20] Naoko Kishi. How does policy focus influence scientific research? Science and Public Policy, 47(1):114– 124, February 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/1/114/5637352.

Kundu:2020:PPI

[KJR20] Oishee Kundu, Andrew D. James, and John Rigby. Public procurement and innovation: a systematic literature review. Science and Public Policy, 47(4):490– 502, August 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/4/490/5903068.

Kang:2021:AEA

[KKK^{+}21] Dong-In Kang, In-Je Kang, Yeong-Ju Kim, Chi Mai Nguyen, and Jae-Yong Choung. Academic entrepreneurship and attentional discrepancy among key stakeholders: Evidence from research universities in Korea. Science and Public Policy, 48(1):1–15, February 2021. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/1/1/6157449.

Kaufmann:2020:HDI

[KKS20] David Kaufmann, Johanna Kuenzler, and Fritz Sager. How (not) to design and implement a large-scale, interdisciplinary research infrastructure. Science and Public Policy, 47(6):818– 828, December 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/6/818/6044344.

Kivimaa:2023:ACR

[KLL23] Paula Kivimaa, Jani Lukkarinen, and David Lazarevic. Analysis of COVID-19 recovery and resilience policy in Finland: a transformative policy mix approach. Science and Public Policy, 50(4):681–694, August 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/4/681/7161512.

Khurana:2023:PNC

[KM23] Mark P. Khurana and Emil Bargmann Madsen. Perspectives on the narrowing and clustering of research trajectories: an epistemic threat to medical progress? Science and Public Policy, 50(3): 559–563, June 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/3/559/6901546.

Karacan:2023:SDG

[KR23] Derya Büyüktanir Karacan and Pierre-Bruno Ruffini. Science diplomacy in the Global South — an introduction. Science and Public Policy, 50(4):742– 748, August 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/50/4/742/7197239.

Kruss:2020:BMM

[Kru20] Glenda Kruss. Balancing multiple mandates: a case study of public research institutes in South Africa. Science and Public Policy, 47(2):149–160, April 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/47/2/149/5634232.

Lagesen:2021:HHD

[Lag21] Vivian Anette Lagesen. How heads of departments find it meaningful to engage with gender balance policies. Science and Public Policy, 48(4):582–591, August 2021. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/48/4/582/6294477.

Lauronen:2020:DUA

[Lau20] Juha-Pekka Lauronen. The dilemmas and uncertainties in assessing the societal impact of research. Science and Public Policy, 47(2): 207–218, April 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/2/207/5700451.

Lawless:2024:BPF

[Law24] Christopher Lawless. Biometrics, presents, futures: the imaginative politics of science–society orderings. Science and Public Policy, 51(2):274–284, April 2024. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/51/2/274/7343994.

Lipp:2023:CCE

[LBCG^{+}23] Benjamin Lipp, Mathieu Baudrin, Carlos Cuevas-Garcia, Federica Pepponi, Magdalena Rozwadowska, and Shelly Tsui. Cocreating end-user roles. Understanding the new variety of user involvement in public sector innovation. Science and Public Policy, 50(1):146–159, February 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/1/146/6754158.

See correction [Ano23b].

Lattu:2023:ILO

[LC23] Annina Lattu and Yuzhuo Cai. Institutional logics in the open science practices of university–industry research collaboration. Science and Public Policy, 50 (5):905–916, October 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/50/5/905/7199923.

Liu:2023:SPF

[LCL23] Pan Liu, Xiaoyan Cui, and Ye Li. Subsidy policies of a fresh supply chain considering the inputs of blockchain traceability service system. Science and Public Policy, 50(1):72– 86, February 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/50/1/72/6702004.

Labib:2022:ETP

[LER^{+}22] Krishma Labib, Natalie Evans, Rea Roje, Panagiotis Kavouras, and Andrea Reyes Elizondo ... Education and training policies for research integrity: Insights from a focus group study. Science and Public Policy, 49(2): 246–266, April 2022. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/49/2/246/6439597.

Lilja:2020:TPA

[Lil20] Erika Lilja. Threat of policy alienation: Exploring the implementation of Open Science policy in research practice. Science and Public Policy, 47(6): 803–817, December 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/6/803/6026552.

Lind:2020:REH

[Lin20] Jonas K. Lind. Resource environment and hierarchy in universities. Science and Public Policy, 47(2): 184–193, April 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/2/184/5652075.

Link:2021:KTF

[Lin21] Albert N. Link. Knowledge transfers from federally funded research and development centers. Science and Public Policy, 48 (4):576–581, August 2021. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/4/576/6292221.

Liu:2022:CSP

[LKZC22] Jing Liu, Xiaoling Kang, Xia Zhang, and Xing Chen. Channels to shape procurement decision-making of public organisations for innovation through framework conditions. Science and Public Policy, 49(3):

474–487, June 2022. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/49/3/474/6521888.

Long:2022:EII

[LL22] Siying Long and Zhongju Liao. Economic incentive instruments and environmental innovation in China: Moderating effect of marketization. Science and Public Policy, 49(4): 553–560, August 2022. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/49/4/553/6540677.

Lee:2022:DBI

[LLM^{+}22] Seonjae Lee, Daehyeon Lim, Youngsu Moon, Hoshin Lee, and Sungjoo Lee. Designing a business intelligence system to support industry analysis and innovation policy. Science and Public Policy, 49(3): 414–426, June 2022. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/49/3/414/6507849.

Li:2023:CAS

[LLRD23] Nancy Li, Markus Luczak-Roesch, and Flavia Donadelli. A computational approach to study the gap and barriers between science and policy. Science and Public Policy, 50(1): 15–29, February 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/1/15/6672664.

Laakso:2023:ESJ

[LM23] Mikael Laakso and Anna-Maija Multas. European scholarly journals from small- and mid-size publishers: mapping journals and public funding mechanisms. Science and Public Policy, 50(3):445– 456, June 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/3/445/7043797.

Loureiro:2020:PSP

[LMR^{+}20] Rafael Loureiro, Omer Markovitch, Crystal S. Riley, H. James Cleaves II, I. Haritina Mogoşanu, Ivan G. Paulino-Lima, Jeffrey J. Marlow, and Sanjoy M. Som. Perspective: Science policy through public engagement. Science and Public Policy, 47(6): 890–891, December 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/6/890/5900807.

Luo:2021:DIN

[LMS21] Junwen Luo, Lai Ma, and Kalpana Shankar. Does the inclusion of non-academic reviewers make any difference for grant impact panels? Science and Public Policy, 48(6):763– 775, December 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/48/6/763/6332840.

Loy:2022:ECC

[Loy22] Loredana Loy. Experts and climate change politicisation. A case study of the Environmental Protection Agency (1983–2015). Science and Public Policy, 49 (5):699–713, October 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/49/5/699/6566070.

Langfeldt:2024:DGP

[LRS24] Liv Langfeldt, Ingvild Reymert, and Silje Marie Svartefoss. Distrust in grant peer review — reasons and remedies. Science and Public Policy, 51 (1):28–41, February 2024. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/51/1/28/7247582.

Link:2020:IDS

[LS20] Albert N. Link and John T.

Scott. Invention disclosures and the slowdown of scientific knowledge. Science and Public Policy, 47(6): 829–833, December 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/6/829/5916437.

Link:2022:AUS

[LSvH22] Albert N. Link, Christopher A. Swann, and Martijn van Hasselt. An assessment of the US Small Business Innovation Research (SBIR) program: a study of project failure. Science and Public Policy, 49(6): 972–978, December 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/6/972/6686585.

Link:2021:PPF

[LW21] Albert N. Link and Caroline S. Wagner. The publicness of publicly funded research. Science and Public Policy, 48(5):757–762, October 2021. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/5/757/6340066.

Liu:2020:DHT

[LXLY20] Huiling Liu, Fei Xing, Bo Li, and Kseniya Yakshtas. Does the hightech enterprise certification policy promote innovation in China? Science and Public Policy, 47(5):678– 688, October 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/47/5/678/6186787.

Ma:2022:DIF

[MA22] Lai Ma and Rachael Agnew. Deconstructing impact: a framework for impact evaluation in grant applications. Science and Public Policy, 49(2):289– 301, April 2022. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/49/2/289/6437551.

Moradi:2023:OSR

[MA23] Sh Moradi and S. Abdi. Open science-related policies in Europe. Science and Public Policy, 50(3): 521–530, June 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/50/3/521/7069603.

See correction [Ano23e].

Marjanovic:2020:IIH

[MAH^{+}20] Sonja Marjanovic, Marlene Altenhofer, Lucy Hocking, Joanna Chataway, and Tom Ling. Innovating for improved healthcare: Sociotechnical and innovation systems perspectives and lessons from the NHS. Science and Public Policy, 47 (2):283–297, April 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/2/283/5722190.

Marteau:2023:ENA

[Mar23] Theresa M. Marteau. Evidenceneglect: addressing a barrier to UK health and climate policy ambitions. Science and Public Policy, 50(6):1103–1109, December 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/6/1103/7202063.

Matus:2020:MAP

[MB20] Kira J. M. Matus and Marie N. Bernal. Media attention and policy response: 21st century chemical regulation in the USA. Science and Public Policy, 47(4):548– 560, August 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/47/4/548/5918728.

Mirnezami:2020:ECT

[MBT20] Seyed Reza Mirnezami, Catherine Beaudry, and Leila Tahmooresnejad. The effect of collaboration with top-funded scholars on scientific production. Science and Public Policy, 47(2): 219–234, April 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/47/2/219/5704800.

MacKillop:2023:MSK

[MCDD23] Eleanor MacKillop, Andrew Connell, James Downe, and Hannah Durrant. Making sense of knowledgebrokering organisations: boundary organisations or policy entrepreneurs? Science and Public Policy, 50(6):950–960, December 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/50/6/950/7204580.

See correction [Ano23d].

Mejdalani:2022:IAI

[MG22] Alexandre Mejdalani and Eduardo Gonçalves. Individual attributes and inventors matching: a study using data from the Brazilian co-patents network.

Science and Public Policy, 49(2):302–312, April 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/2/302/6440317.

Marais:2021:SFF

[MGM^{+}21] Annica Marais, Sara S. (Saartjie)

Grobbelaar, Isabel Meyer, Denzil Kennon, and Marlien Herselman. Supporting the formation and functioning of innovation platforms in healthcare value chains. Science and Public Policy, 48(1):105–121, February 2021. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/1/105/6026554.

Mo:2023:EPR

[Mo23] Jung Youn Mo. Environmental policy and R&D productivity: a case study from the Korean Emissions Trading Scheme. Science and Public Policy, 50(1): 120–128, February 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/1/120/6705400.

Moujaes:2024:MSS

[Mou24] Ghinwa Moujaes. Moving to Smart Specialization for sustainability: the implications on the design of monitoring indicators. Science and Public Policy, 51(1): 127–143, February 2024. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/51/1/127/7308807.

Morisson:2023:PEE

[MP23] Arnault Morisson and Evangelia Petridou. Policy entrepreneurs, entrepreneurial strategies, and institutional contexts in Interreg Europe. Science and Public Policy, 50 (4):579–590, August 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/50/4/579/7153042.

Moncada-Paterno-Castello:2020:CRI

[MPCAC20] Pietro Moncada-Paternò-Castello, Sara Amoroso, and Michele Cincera. Corporate R&D intensity decomposition: different data, different results? Science and Public Policy, 47 (4):458–473, August 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/47/4/458/5897678.

Minssen:2020:CTD

[MRB20] Timo Minssen, Neethu Rajam, and Marcel Bogers. Clinical trial data transparency and GDPR compliance: Implications for data sharing and open innovation. Science and Public Policy, 47(5):616– 626, October 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/5/616/5780440.

Machacek:2022:RIM

[MSF^{+}22] Vít Macháček, Martin Srholec, Márcia R Ferreira, Nicolas Robinson-Garcia, and Rodrigo Costas. Researchers’ institutional mobility: bibliometric evidence on academic inbreeding and internationalization. Science and Public Policy, 49(1):85– 97, February 2022. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/1/85/6368540.

Mora-Valentin:2022:ASU

[MVNSdUC22] Eva-María Mora-Valentín, Juan-José Nájera-Sánchez, and Marta Ortiz de Urbina-Criado. Assessment of success in university–industry cooperation literature: a bibliographic coupling analysis. Science and Public Policy, 49(5):686–698, October 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/49/5/686/6564065.

Marini:2021:GBC

[MY21] Giulio Marini and Lili Yang. Globally bred Chinese talents returning home: an analysis of a reverse brain-drain flagship policy. Science and Public Policy, 48(4):541– 552, August 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/48/4/541/6290435.

NogueiradePaivaBritto:2020:PCK

[NCTd20] Jorge Nogueira de Paiva Britto, Leonardo Costa Ribeiro, Lucas Teixeira Araújo, and Eduardo da Motta e Albuquerque. Patent citations, knowledge flows, and catching-up: Evidences of different national experiences for the period 1982–2006. Science and Public Policy, 47(6):788– 802, December 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/47/6/788/6025446.

Nelson:2024:MDS

[Nel24]  John P. Nelson.  The

micro-dynamics of scientific choice: research project motivations among public affairs academics. Science and Public Policy, 51(1):

149–161, February 2024. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/51/1/149/7308801.

Nikolova:2024:RRE

[Nik24] Blagovesta I. Nikolova.

From responsibility to risk: ethics in the Bermuda Triangle of EU research and innovation policy. Science and Public Policy, 51(2): 207–217, April 2024. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/51/2/207/7329135.

Nylen:2023:MOI

[NJV23] Erkki-Jussi Nylén, Jan-Erik Johanson, and Jarmo Vakkuri. Mission-oriented innovation policy as a hybridisation process: the case of transforming a national fertilising system. Science and Public Policy, 50(3):509–520, June 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/3/509/7059653.

Nawaz:2021:DLS

[NK21] Sara Nawaz and Milind Kandlikar. Drawing lines in the sand? Paths forward for triggering regulation of gene-edited crops. Science and Public Policy, 48(2):246–256, April 2021. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/48/2/246/6173809.

Nienow:2024:BTM

[NLOL24] Sara Nienow, Olena Leonchuk, Alan C. O’Connor, and Albert N. Link. Bringing technology to market: National Heart, Lung, and Blood Institute SBIR Phase IIB projects. Science and Public Policy, 51(1): 144–148, February 2024. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/51/1/144/7308804.

Ng:2023:RIN

[NLP23] Huey Yuen Ng, Yining Luo, and Hyunkyu Park. The role of intermediaries in nurturing innovation ecosystems: a case study of Singapore’s manufacturing sector. Science and Public Policy, 50(3): 382–397, June 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/50/3/382/6960882.

Nelhans:2022:ATW

[NN22] Gustaf Nelhans and Jan Nolin. Anti-transparency within the EU shift to open science. Science and Public Policy, 49(6):928– 941, December 2022. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/6/928/6730548.

Nishimura:2022:MUI

[NNAH22] Junichi Nishimura, Sadao Nagaoka, Shinichi Akaike, and Mitsuaki Hosono. Making university and industry research collaboration: Evidence from coinventions in Japan. Science and Public Policy, 49 (2):268–288, April 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/2/268/6462107.

Nasution:2022:EBC

[NNS^{+}22] Bahrul Ilmi Nasution, Yudhistira Nugraha, Andi Sulasikin, Hansen Wiguna, Juan Intan Kanggrawan, Alex Lukmanto Suherman, Ngabila Salama, and Dwi Oktavia. An evidencebased culture: COVID-19 positivity factors during the asymptomatic occurrence in Jakarta, Indonesia. Science and Public Policy, 49(1):115–126, February 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/1/115/6395853.

Nascia:2021:SLP

[NPZ21] Leopoldo Nascia, Mario Pianta, and Thomas Zacharewicz. Staying or leaving? Patterns and determinants of Italian researchers’ migration. Science and Public Policy, 48(2):200–211, April 2021. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/48/2/200/6134320.

Novella:2023:WNT

[NRSA23] Rafael Novella, David Rosas-Shady, and Alfredo Alvarado. Are we nearly there yet? New technology adoption and labor demand in Peru. Science and Public Policy, 50(4):565– 578, August 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/50/4/565/7084881.

Olson:2020:KPS

[OdS20] Julia Olson and Patricia Pinto da Silva. Knowledge production at the science–policy interface: Lessons from fisheries scientists. Science and Public Policy, 47(1):47–55, February 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/1/47/5614237.

Otten:2022:ABI

[OF22] Caitlin Drummond Otten and Baruch Fischhoff. Assessing broader impacts of funded research: the US National Science Foundation v. Lamar Smith. Science and Public Policy, 49 (2):313–323, April 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/2/313/6461024.

Onoda:2022:BSR

[OI22] Takashi Onoda and Yasunobu Ito. On the boundary of services and research collaborations in Japanese state-of-the-art academic research infrastructures. Science and Public Policy, 49(3):488–498, June 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/3/488/6526952.

Okamuro:2021:EMP

[ON21] Hiroyuki Okamuro and Junichi Nishimura. Effects of multilevel policy mix of public R&D subsidies: Empirical evidence from Japanese local SMEs. Science and Public Policy, 48(6):829– 840, December 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/48/6/829/6353940.

Orjuela-Ramirez:2024:DEI

[ORZJU24] Guillermo Orjuela-Ramirez, Julio Cesar Zuluaga-Jimenez, and David Urbano. Drivers of eco-innovation: the role of appropriability strategies and complementary assets. Science and Public Policy, 51(2):309–323, April 2024. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/51/2/309/7421929.

OShaughnessy:2023:WGA

[OSV^{+}23] Matthew R. O’Shaughnessy, Daniel S. Schiff, Lav R. Varshney, Christopher J. Rozell, and Mark A. Davenport. What governs attitudes toward artificial intelligence adoption and governance? Science and Public Policy, 50 (2):161–176, April 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

                                        https://academic.oup.
                                        com/spp/article/50/2/
                                        161/6760880.

Pang:2020:HDI

[Pan20]  Zhenjing Pang.  How

does information dependence affect public attitudes towards genetically modified technology in China? Science and Public Policy, 47(1):125– 136, February 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/1/125/5644095.

Parthasarathy:2020:AIP

[Par20] Shobita Parthasarathy. Approaching intellectual property scholarship differently: a qualitative research review and agenda. Science and Public Policy, 47 (5):627–637, October 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/5/627/5762653.

Piro:2020:UCP

[PBSA20] Fredrik Niclas Piro, Pål Børing, Lisa Scordato, and Dag W. Aksnes. University characteristics and probabilities for funding of proposals in the European Framework Programs. Science and Public Policy, 47 (4):581–593, August 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/4/581/5920444.

Popa:2020:PAF

[PBW20] Eugen Octav Popa, Vincent Blok, and Renate Wesselink. A processual approach to friction in quadruple helix collaborations. Science and Public Policy, 47(6):876–889, December 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/47/6/876/6025445.

Petrin:2023:CBI

[PDM23] Renata Petrin, Roberto Gonzalez Duarte, and Irene Kazumi Miura. Coevolution between institutions and scientific organizations: The case of IMPA. Science and Public Policy, 50 (2):287–303, April 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

                                        https://academic.oup.
                                        com/spp/article/50/2/
                                        287/6855248.

Polonen:2020:AIP

[PEG20] Janne Pölönen, Tim Engels, and Raf Guns. Ambiguity in identification of peer-reviewed publications in the Finnish and Flemish performance-based research funding systems. Science and Public Policy, 47(1):1–15, February 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/47/1/1/5564001.

Peng:2023:LMC

[Pen23] Huidong Peng. Local market, central government support, and local governments’ homegrown development strategy in hightech industries. Science and Public Policy, 50(6):1073– 1090, December 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/6/1073/7240641.

Pesch:2021:IIT

[Pes21] Udo Pesch. Imaginaries of innovation: Turning technology development into a public issue. Science and Public Policy, 48(2): 257–264, April 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/2/257/6169404.

Pozzo:2020:WDC

[PFPV20] Riccardo Pozzo, Andrea Filippetti, Mario Paolucci, and Vania Virgili. What does cultural innovation stand for? Dimensions, processes, outcomes of a new innovation category. Science and Public Policy, 47(3):425–433, June 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/3/425/5830820.

Penna:2023:EIM

[PGM23] Caetano C. R. Penna, Oscar Yandy Romero Goyeneche, and Cristian Matti. Exploring indicators for monitoring sociotechnical system transitions through portfolio networks. Science and Public Policy, 50(4):719– 741, August 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/50/4/719/7189759.

Peterson:2023:OPW

[PH23] Helen Peterson and Liisa Husu. Online panel work through a gender lens: implications of digital peer review meetings. Science and Public Policy, 50(3): 371–381, June 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/50/3/371/6958163.

Philipps:2022:RFR

[Phi22] Axel Philipps. Research funding randomly allocated? A survey of scientists’ views on peer review and lottery. Science and Public Policy, 49(3): 365–377, June 2022. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/49/3/365/6463570.

Parker:2022:RUT

[PL22] Rachel Parker and Petra Lundgren. The role of universities in transformative innovation policy. Science and Public Policy, 49(2): 159–167, April 2022. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/2/159/6430222.

Pillai:2020:TRI

[PM20] Vishnu Sivarudran Pillai and Kira J. M. Matus. Towards a responsible integration of artificial intelligence technology in the construction sector. Science and Public Policy, 47 (5):689–704, October 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/5/689/6104655.

Polejack:2023:CSD

[Pol23] Andrei Polejack. Coloniality in science diplomacy — evidence from the Atlantic Ocean. Science and Public Policy, 50(4):759– 770, August 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/4/759/7187282.

Prainsack:2020:MEO

[Pra20] Barbara Prainsack. The meaning and enactment of openness in Personalised and Precision Medicine.

Science and Public Policy, 47(5):647–654, October 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/47/5/647/5760336.

Peltola:2023:RRC

[PSK^{+}23] Taru Peltola, Sanna-Riikka Saarela, Juha M. Kotilainen, Tapio Litmanen, Jani Lukkarinen, Ismo Pölönen, Outi Ratamäki, Heli Saarikoski, Miikka Salo, and Suvi Vikström. Researcher roles in collaborative governance interventions. Science and Public Policy, 50(5):871– 880, October 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/50/5/871/7197920.

Pezzoni:2024:GBT

[PV24] Michele Pezzoni and Fabiana Visentin. Gender bias in team formation: the case of the European Science Foundation’s grants. Science and Public Policy, 51(2):247–260, April 2024. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/51/2/247/7332178.

Qian:2023:GSP

[Qia23] Furong Qian. Government subsidies, private R&D, and global value chains position: the moderating role of task complexity. Science and Public Policy, 50(6): 988–1000, December 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/6/988/7469002.

Quintero:2023:DEU

[QS23] Natalia Dévora Quintero and Gonzalo León Serrano. Discussion on the entrepreneurial university in Spain: The case of the Madrid Region. Science and Public Policy, 50 (2):253–274, April 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

https://academic.oup.
com/spp/article/50/2/
253/6827976.

Robinson:2023:GSD

[RAB^{+}23] Sam Robinson, Matthew Adamson, Gordon Barrett, Lif Lund Jacobsen, Simone Turchetti, Aya Homei, Péter Marton, Leah Aronowsky, Iqra Choudry, Johan Gärdebo, Jaehwan Hyun, Gerardo Ienna, Carringtone Kinyanjui, Beatriz Martínez-Rius, Júlia Mascarello, Doubravka Olsakova, Giulia Rispoli, and Waqar Zaidi. The globalization of science diplomacy in the early 1970s: a historical exploration. Science and Public Policy, 50 (4):749–758, August 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/50/4/749/7193006.

Robinson:2021:EFQ

[RAJV21] Douglas K. R. Robinson, Marc Audétat, Pierre-Benoit Joly, and Harro Van Lente. Enemies of the future? Questioning the regimes of promising in emerging science and technology. Science and Public Policy, 48(6):814– 817, December 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/48/6/814/6338502.

Razack:2022:CTT

[RAM22] Habeeb Ibrahim Abdul Razack, Jesil Mathew Aranjani, and Sam T. Mathew. Clinical trial transparency regulations: Implications to various scholarly publishing stakeholders. Science and Public Policy, 49(6):951–961, December 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/49/6/951/6654535.

Rohracher:2023:MIL

[RCK23] Harald Rohracher, Lars Coenen, and Olga Kordas. Mission incomplete: Layered practices of monitoring and evaluation in Swedish transformative innovation policy. Science and Public Policy, 50(2):336–349, April 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

https://academic.oup.
com/spp/article/50/2/
336/6889782.

Reale:2022:FES

[Rea22]  Emanuela Reale.  Fac-

tors enabling social impact: The importance of institutional entrepreneurship in social science research. Science and Public Policy, 49 (4):632–642, August 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/4/632/6551528.

Reichenbach:2022:OPR

[REL22] Rebecca Reichenbach, Christoph Eberl, and Jörg Lindenmeier. Online platforms for research data: a requirements and cost analysis. Science and Public Policy, 49(4):598–608, August 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/49/4/598/6554496.

Ruffini:2023:SDN

[RK23] Pierre-Bruno Ruffini and Olga Krasnyak. Science diplomacy from a nationstate’s perspective: a general framing and its application to Global South countries. Science and Public Policy, 50(4):771– 781, August 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/50/4/771/7193007.

Radosevic:2023:EAT

[RKT23] Slavo Radosevic, Despina Kanellou, and George Tsekouras. The experimentation– accountability trade-off in innovation and industrial policy: are learning networks the solution? Science and Public Policy, 50 (4):655–669, August 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/50/4/655/7153046.

Rainey:2021:PNC

[RML^{+}21] Stephen Rainey, Maru Mormina, Sapfo Lignou, Joseph Nguyen, and Paula Larsson. The post-normal challenges of COVID-19: Constructing effective and legitimate responses. Science and Public Policy, 48 (4):592–601, August 2021. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/4/592/6294463.

Ruess:2023:ORT

[RMP23] Anja K. Ruess, Ruth Müller, and Sebastian M. Pfotenhauer. Opportunity or responsibility? Tracing co-creation in the European policy discourse. Science and Public Policy, 50(3): 433–444, June 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/3/433/7044673.

Rohracher:2024:NME

[RO24] Harald Rohracher and Michael Ornetzeder. Navigating missions: experiences from a long-term R&I programme to transform the building sector in Austria. Science and Public Policy, 51(1):67– 79, February 2024. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/51/1/67/7281957.

Rodder:2023:UBS

[RP23] Simone Rödder and Christopher Niklas Pavenstädt. ‘Unite behind the science!’ Climate movements’ use of scientific evidence in narratives on socio-ecological futures. Science and Public Policy, 50(1):30–41, February 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/1/30/6674005.

Rossmann:2020:KNM

[RR20] Maximilian Roßmann and Christine Rösch. Keynarratives of microalgae nutrition: Exploring futures through a public Delphi survey in Germany. Science and Public Policy, 47(1):137–147, February 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/1/137/5645159.

Ruland:2023:SDG

[RRC^{+}23] Anna-Lena Rüland, Nicolas Rüffin, Katharina Cramer, Prosper Ngabonziza, Manoj Saxena, and Stefan Skupien. Science diplomacy from the Global South: the case of intergovernmental science organizations. Science and Public Policy, 50(4):782– 793, August 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/4/782/7197178.

Rudek:2022:CIS

[Rud22] Tadeusz Józef Rudek. Capturing the invisible. Sociotechnical imaginaries of energy. The critical overview. Science and Public Policy, 49(2):219– 245, April 2022. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/2/219/6429333.

Rzymski:2022:ROW

[Rzy22] Piotr Rzymski. Retraction in the online world — shall we rethink the policy? Science and Public Policy, 49(3):378–381, June 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/3/378/6459717.

Scordato:2022:WTI

[SBH^{+}22] Lisa Scordato, Markus M.

Bugge, Teis Hansen, Anne Tanner, and Olav Wicken. Walking the talk? Innovation policy approaches to unleash the transformative potentials of the Nordic bioeconomy. Science and Public Policy, 49(2):324– 346, April 2022. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/2/324/6447025.

Salomaa:2021:UTM

[SC21a] Maria Salomaa and David Charles. The university third mission and the European structural funds in peripheral regions: Insights from Finland. Science and Public Policy, 48(3): 352–363, June 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/48/3/352/6126876.

Stilgoe:2021:RAL

[SC21b] Jack Stilgoe and Tom Cohen. Rejecting acceptance: learning from public dialogue on self-driving vehicles. Science and Public Policy, 48(6):849–859, December 2021. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/48/6/849/6355299.

Sharif:2022:CAI

[SC22] Naubahar Sharif and Kevin Chandra. A comparative analysis of innovation policies in Hong Kong and Shenzhen within the Greater Bay Area initiative. Science and Public Policy, 49(1):54–71, February 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/1/54/6366459.

Son:2021:ABP

[SCY21] Hosung Son, Yanghon Chung, and Sangpil Yoon. Is the alignment between public research organisations’ R&D competence and policies really critical for technology transfer? Science and Public Policy, 48(1):93–104, February 2021. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/1/93/6026553.

Santos:2023:UIC

[SDM23] Telmo N. Santos, José G.

Dias, and Sandro Mendonça. University–industry cooperation: a taxonomy of intermediaries. Science and Public Policy, 50(3): 457–490, June 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/3/457/7043715.

Suarez:2023:OUC

[SFP23] Diana Suarez, Florencia Fiorentin, and Mariano Pereira. Observable and unobservable causes of the gender gap in S&T funding for young researchers. Science and Public Policy, 50 (4):579–590, August 2023.

CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/4/579/7084754.

Schaeffer:2024:WME

[SFQdM24] Paola Rücker Schaeffer, Bruno Brandão Fischer, Sérgio Queiroz, and Gustavo Hermínio Salati Marcondes de Moraes. What makes an entrepreneurial university? Institutional moderators of ecosystem impacts in a developing country. Science and Public Policy, 51(1):108–126, February 2024. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/51/1/108/7308805.

Scholten:2021:FFA

[SFvD^{+}21] Wout Scholten, Thomas P. Franssen, Leonie van Drooge, Sarah de Rijcke, and Laurens K. Hessels. Funding for few, anticipation among all: Effects of excellence funding on academic research groups. Science and Public Policy, 48(2): 265–275, April 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/2/265/6184850.

Shi:2023:MIP

[SGDW23] Xing Shi, Yating Guo, Huiping Dong, and Shuai Wang. Multilevel innovation policy mix in China: do local programmes complement national programmes? Science and Public Policy, 50(6):1029– 1043, December 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/6/1029/7236567.

Sefyrin:2021:ADD

[SGW21] Johanna Sefyrin, Mariana Gustafsson, and Elin Wihlborg. Addressing digital diversity: Care matters in vulnerable digital relations in a Swedish library context. Science and Public Policy, 48(6):841– 848, December 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/6/841/6353893.

Sarkki:2020:HBO

[SHK^{+}20] Simo Sarkki, Hannu I.

Heikkinen, Teresa Komu, Mari Partanen, Karoliina Vanhanen, and Élise Lépy. How boundary objects help to perform roles of science arbiter, honest broker, and issue advocate. Science and Public Policy, 47(2): 161–171, April 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/47/2/161/5647310.

Simons:2020:HPB

[SHRA20] Arno Simons, Barbara Hendriks, Martin Reinhart, and Faten Ahmed. How practitioners between bench and bedside evaluate biomedical translation? Science and Public Policy, 47(4):561–570, August 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/47/4/561/5896073.

Sideri:2020:PPD

[Sid20] Katerina Sideri. Prospect patents, data markets, and the commons in datadriven medicine: Openness and the political economy of intellectual property rights. Science and Public Policy, 47(5):668– 677, October 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/47/5/668/5809099.

Siler:2024:GLM

[Sil24] Kyle Siler. Gerontocracy, labor market bottlenecks, and generational crises in modern science. Science and Public Policy, 51(2): 179–191, April 2024. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/51/2/179/7323635.

Shekar:2022:DII

[SJ22] K. Chandra Shekar and K. J. Joseph. Determinants of innovation and interactive learning in informal manufacturing enterprises in India. Science and Public Policy, 49(3): 427–440, June 2022. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/3/427/6554494.

Svingen:2024:MSC

[SJ24] Marit Svingen and Lisbeth Jahren. Making space for CRISPR: scientists’ translation work to make gene editing a legitimate technology. Science and Public Policy, 51(1): 15–27, February 2024. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/51/1/15/7247522.

Sjoo:2023:GMR

[SK23] Karolin Sjöö and Wolfgang Kaltenbrunner. Gender mainstreaming research funding: a study of effects on STEM research proposals. Science and Public Policy, 50(2):304–317, April 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

https://academic.oup.
com/spp/article/50/2/
304/6854962.

Sachini:2021:RRC

[SLSC^{+}21] Evi Sachini, Lois Labrianidis, Konstantinos Sioumalas-Christodoulou, Charalampos Chrysomallidis, Galatios Siganos, Agathi Belouli, and Nikolaos Karampekios. Research on researchers. Coping during COVID-19. Results on a nation-wide survey. Science and Public Policy, 48 (4):451–461, August 2021. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/4/451/6213373.

Spano:2024:LID

[SMPB24] Eduardo Spanó, Rafael Monnerat, Carlos Américo Pacheco, and Maria Beatriz Machado Bonacelli. Legal–institutional design and dynamic capabilities for mission-oriented innovation agencies: a new framework. Science and Public Policy, 51(2):192– 206, April 2024. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/51/2/192/7329134.

Solbu:2021:FBC

[Sol21] Gisle Solbu. Frictions in the bioeconomy? A case study of policy translations and innovation practices. Science and Public Policy, 48(6):911–920, December 2021. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/6/911/6409637.

Scholvinck:2020:PIA

[SPB20] Anne-Floor M. Schölvinck, Carina A. C. M. Pittens, and Jacqueline E. W. Broerse. Patient involvement in agenda-setting processes in health research policy: a boundary work perspective. Science and Public Policy, 47(2):246– 255, April 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/2/246/5766086.

Schmidt:2023:PME

[SPB23] Evanthia Kalpazidou Schmidt, Rachel Palmén, and Susanne Bührer. Policymaking and evaluation of gender equality programmes: Context, power, and resistance in the transformation process. Science and Public Policy, 50 (2):206–218, April 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

                                        https://academic.oup.
                                        com/spp/article/50/2/
                                        206/6795957.

Sondergaard:2021:VPC

[ST21] Helle Alsted Søndergaard and John Thøgersen. The value priorities of consumer innovators. Science and Public Policy, 48(5):679– 696, October 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/48/5/679/6324092.

Sormani:2022:APM

[SU22] E. Sormani and K. Uude. Academics’ prosocial motivation for engagement with society: The case of German academics in health science. Science and Public Policy, 49(6):962–971, December 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/49/6/962/6657698.

Stage:2023:FLF

[SU23] Andreas Kjær Stage and Ea Høg Utoft. Fun and less fun funding: the experiential affordances of research grant conditions. Science and Public Policy, 50(6): 1091–1102, December 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/6/1091/7251151.

Scott:2021:HTI

[SWA^{+}21] Troy Scott, Amanda Walsh, Benjamin Anderson, Alan O’Connor, and Gregory Tassey. High-tech infrastructure and economic growth: The Materials Genome Initiative. Science and Public Policy, 48 (5):649–661, October 2021. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/5/649/6311311.

Schultz:2024:PPS

[SWAD24] Émilien Schultz, Jeremy K.

Ward, and Laëtitia Atlani-Duault. Public perception of scientific advisory bodies: the case of France’s Covid-19 Scientific Council. Science and Public Policy, 51(2):236–246, April 2024. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/51/2/236/7330826.

Sharif:2019:RGC

[SX19] Naubahar Sharif and Jack Linzhou Xing. Restricted generalizability of city innovation policies: The case of ehailing in China. Science and Public Policy, 46(6): 805–819, December 2019.

CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/46/6/805/5564003.

See erratum [SX20].

Sharif:2020:ERG

[SX20] Naubahar Sharif and Jack Linzhou Xing. Erratum: Restricted generalizability of city innovation policies: The case of e-hailing in China. Science and Public Policy, 47(1):148, February 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/47/1/148/5592229.

See [SX19].

Tassey:2024:FAT

[Tas24]  Gregory Tassey.  A

four-asset technology-based growth policy. Science and Public Policy, 51(1): 42–54, February 2024. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/51/1/42/7255510.

Tahmooresnejad:2021:SNE

[TBM21] Leila Tahmooresnejad, Catherine Beaudry, and Seyed Reza Mirnezami. The study of network effects on research impact in Africa. Science and Public Policy, 48(4): 462–473, August 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/4/462/6215147.

Tang:2021:DSE

[TCWZ21] Li Tang, Cong Cao, Zheng Wang, and Zhuo Zhou. Decoupling in science and education: a collateral damage beyond deteriorating US–China relations. Science and Public Policy, 48 (5):630–634, October 2021. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/5/630/6322422.

Terbullino:2023:CEL

[Ter23] Pavel Gabriel Corilloclla Terbullino. Centres of excellence in Latin America: how do these differ from other experiences? Science and Public Policy, 50 (4):591–602, August 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/4/591/7095646.

Tellmann:2022:OSB

[TG22] Silje Maria Tellmann and Magnus Gulbrandsen. The other side of the boundary: Productive interactions seen from the policy side. Science and Public Policy, 49(4):621– 631, August 2022. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/49/4/621/6551529.

Tang:2020:IDI

[THPT20] Yongli Tang, Xinyue Hu, Claudio Petti, and Matthias Thürer. Institution-driven innovation in Guangdong firms: Moderating effects of in-house formal R&D and industrial environment turbulence. Science and Public Policy, 47(2):194– 206, April 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/47/2/194/5684867.

Takano:2021:LRS

[TO21] Keisuke Takano and Hiroyuki Okamuro. Local R&D support as a driver of network diversification: a cross-regional comparison in Japan. Science and Public Policy, 48(6):776– 787, December 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/48/6/776/6333642.

Tatsch:2022:KNB

[TRdRABS22] Ana Lúcia Tatsch, Janaina Ruffoni, Marisa dos Reis A Botelho, and Rafael Stefani. Knowledge networks in Brazil’s health sciences. Science and Public Policy, 49(1):72– 84, February 2022. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/1/72/6359621.

Teirlinck:2023:IRI

[TS23] Peter Teirlinck and André Spithoven. Improving the Regional Innovation Scoreboard for policy: how about innovation efficiency? Science and Public Policy, 50(6):1001– 1017, December 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/6/1001/7233579.

Terzi:2023:EIP

[TSS23] Alessio Terzi, Monika Sherwood, and Aneil Singh. European industrial policy for the green and digital revolution. Science and Public Policy, 50(5):842– 857, October 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/5/842/7192958.

Todtling:2020:POG

[TTF20] Franz Tödtling, Michaela Trippl, and Alexandra Frangenheim. Policy options for green regional development: Adopting a production and application perspective. Science and Public Policy, 47(6):865– 875, December 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/47/6/865/6026550.

Vienni-Baptista:2022:EHW

[VBFLP22] Bianca Vienni-Baptista, Isabel Fletcher, Catherine Lyall, and Christian Pohl. Embracing heterogeneity: Why plural understandings strengthen interdisciplinarity and transdisciplinarity. Science and Public Policy, 49(6):865– 877, December 2022. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/49/6/865/6619578.

See correction [Ano23c].

Verwoerd:2023:NSK

[VBK^{+}23] Lisa Verwoerd, Hilde Brouwers, Eva Kunseler, Barbara Regeer, and Evelien de Hoop. Negotiating space for knowledge coproduction. Science and Public Policy, 50(1):59– 71, February 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/50/1/59/6691206.

vandeKlippe:2023:PRP

[vdKYYWR23] Wouter van de Klippe, Alfredo Yegros-Yegros, Tim Willemse, and Ismael Rafols. Priorities in research portfolios: exploring the need for upstream research in cardiometabolic and mental health. Science and Public Policy, 50(6): 961–976, December 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/6/961/7204582.

Vincenzi:2022:WDB

[Vin22] Marco Vincenzi. Why do Big Science projects exist? The role of social preferences. Science and Public Policy, 49(6):853– 864, December 2022. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/6/853/6618501.

Vega-Jurado:2020:VCE

[VJMHdLNA20] Jaider Vega-Jurado, Liney Manjarrés-Henríquez, Ignacio Fernández de Lucio, and Gloria Naranjo-Africano. A virtuous circle? The effects of university– industry relationships in a region with low absorptive capacity. Science and Public Policy, 47(4):503– 513, August 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/47/4/503/5896080.

Varumo:2020:CSD

[VPK20] Liisa Varumo, Riikka Paloniemi, and Eszter Kelemen. Challenges and solutions in developing legitimate online participation for EU biodiversity and ecosystem services policies. Science and Public Policy, 47(4):571– 580, August 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/47/4/571/5989932.

Vicente-Saez:2021:OSS

[VSGMF21] Ruben Vicente-Saez, Robin Gustafsson, and Clara Martinez-Fuentes. Opening up science for a sustainable world: an expansive normative structure of open science in the digital era. Science and Public Policy, 48(6):799– 813, December 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/48/6/799/6340010.

Vehvilainen:2022:SGE

[VV22] Marja Vehviläinen and Liekki Valaskivi. Situated gender equality in regional research and innovation: Collaborative knowledge production. Science and Public Policy, 49(4): 561–572, August 2022. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/4/561/6540130.

vanWelie:2020:ISF

[vWBT20] Mara J. van Welie, Wouter P. C. Boon, and Bernhard Truffer. Innovation system formation in international development cooperation: The role of intermediaries in urban sanitation. Science and Public Policy, 47(3):333–347, June 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/3/333/5740778.

Woodson:2022:IWA

[WB22] Thomas Woodson and Sophia Boutilier. Impacts for whom? Assessing inequalities in NSF-funded broader impacts using the inclusion-immediacy criterion. Science and Public Policy, 49(2):168–178, April 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/2/168/6430224.

West:2022:PDR

[WBM22] William F. West, Leah W.

Buchman, and Raul F.

Medina. Public deliberation and the regulation of gene drive in the USA. Science and Public Policy, 49(6):843–852, December 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/6/843/6618762.

Whetsell:2021:DCS

[WDJ^{+}21] Travis A. Whetsell, Ana-Maria Dimand, Koen Jonkers, Jeroen Baas, and Caroline S. Wagner. Democracy, complexity, and science: Exploring structural sources of national scientific performance. Science and Public Policy, 48(5):697– 711, October 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/5/697/6325624.

Wang:2024:LQC

[WHH24] Jing Wang, Willem Halffman, and Serge P. J. M. Horbach. Listing quality: Chinese journal lists in incoherent valuation regimes. Science and Public Policy, 51(1):55– 66, February 2024. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/51/1/55/7279195.

Wittmann:2021:GVM

[WHL^{+}21] Florian Wittmann, Miriam Hufnagl, Ralf Lindner, Florian Roth, and Jakob Edler. Governing varieties of mission-oriented innovation policies: a new typology. Science and Public Policy, 48(5):727– 738, October 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/5/727/6332628.

Williams:2022:WCM

[Wil22]  Kate Williams.  What

counts: Making sense of metrics of research value. Science and Public Policy, 49(3):518–531, June 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/3/518/6550865.

Wang:2020:EIP

[WJZ20] Xiaozhen Wang, Zihao Jiang, and Ying Zheng. Effect of innovation policy mix on innovation efficiency: Evidence from Chinese wind power industry chain. Science and Public Policy, 47(1):31– 46, February 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/1/31/5564329.

Whetsell:2020:BPP

[WLW20] Travis A. Whetsell, Michael J. Leiblein, and Caroline S. Wagner. Between promise and performance: Science and technology policy implementation through network governance. Science and Public Policy, 47(1): 78–91, February 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/47/1/78/5618663.

Woelert:2021:KSC

[Woe21] Peter Woelert. ‘Key selection criteria’ as administrative devices: an investigation of academic bureaucratization at Australian universities. Science and Public Policy, 48(1):27– 36, February 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/48/1/27/6026547.

Wienroth:2021:PER

[WS21] Matthias Wienroth and Jackie Leach Scully. Promissory ethical regimes: publics and public goods in genome editing for human health. Science and Public Policy, 48(6):788–798, December 2021. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/6/788/6338190.

Wagner:2023:CUF

[WS23] Caroline S. Wagner and Denis F. Simon. China’s use of formal science and technology agreements as a tool of diplomacy. Science and Public Policy, 50(4):807– 817, August 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/4/807/7209674.

Weingart:2022:DEL

[WvSG22] Peter Weingart, François van Schalkwyk, and Lars Guenther. Democratic and expert legitimacy: Science, politics and the public during the COVID-19 pandemic. Science and Public Policy, 49(3):499– 517, June 2022. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/3/499/6526989.

Wanzenbock:2020:FMO

[WWF^{+}20] Iris Wanzenböck, Joeri H.

Wesseling, Koen Frenken, Marko P. Hekkert, and K. Matthias Weber. A framework for missionoriented innovation policy: Alternative pathways through the problem– solution space. Science and Public Policy, 47(4):474– 489, August 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/47/4/474/5874664.

Yeh:2022:WCK

[wY22] Chih wei Yeh. Working with curiosity: Knowledge transfer practitioners’ ambivalence at CERN. Science and Public Policy, 49 (5):781–790, October 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/49/5/781/6611714.

See correction [Ano22].

Wu:2023:DOP

[WZ23] Ying Wu and Yi Zhang. Does offensive patent insurance promote corporate innovation? Evidence from a quasi-experiment. Science and Public Policy, 50 (2):275–286, April 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

                                        https://academic.oup.
                                        com/spp/article/50/2/
                                        275/6845580.

Xu:2023:CIV

[XG23] Nini Xu and Junhua Guo. Can innovation vouchers promote SMEs’ innovation in China: a flexible conditional differencein-differences analysis. Science and Public Policy, 50 (2):230–242, April 2023. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL

https://academic.oup.
com/spp/article/50/2/
230/6825384.

Xie:2022:ICA

[XHQ^{+}22] Ruiyao Xie, Chaocheng He, Shan Qiao, Xiaoming Li, Jiang Wu, Weiming Tang, Wentian Cui, and Qingpeng Zhang. The impacts of centers for AIDS research program and its enlargement on HIV/AIDS research collaboration. Science and Public Policy, 49 (5):714–727, October 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/5/714/6584493.

Xia:2022:EED

[XLL22] Qing Xia, Lanjian Liu, and Xiaoping Li. Evolution of the emergency drug innovation network and policy implications: Evidence from COVID-19 drug patents. Science and Public Policy, 49(3):403–413, June 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/3/403/6507851.

Younes:2020:CII

[YAB^{+}20] George Abi Younes, Charles Ayoubi, Omar Ballester, Gabriele Cristelli, Gaétan de Rassenfosse, Dominique Foray, Patrick Gaulé, Gabriele Pellegrino, Matthias van den Heuvel, Elizabeth Webster, and Ling Zhou. COVID-19: Insights from innovation economists. Science and Public Policy, 47 (5):733–745, October 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/47/5/733/5865508.

Yang:2021:HDL

[Yan21] Lihua Yang. How does the local knowledge of scientists influence their impact on governance performance? Science and Public Policy, 48(3):334– 352, June 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/48/3/334/6158534.

Yang:2022:PEY

[YCL22] Xi Yang, Xinlan Cai, and Tingsong Li. Peer effects of the young returnee scientists: Evidence from the state key laboratories in China. Science and Public Policy, 49(5):739– 750, October 2022. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/5/739/6583384.

Yeh:2021:HID

[Yeh21] Chih-Wei Yeh. A hierarchy of interests: Discursive practices on the value of particle and highenergy physics. Science and Public Policy, 48(3): 325–333, June 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/3/325/6291409.

Yoon:2021:DER

[YH21] Haeyeon Yoon and Almas Heshmati. Do environmental regulations affect FDI decisions? The pollution haven hypothesis revisited. Science and Public Policy, 48(1):122– 131, February 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/1/122/5922858.

Youtie:2021:ICA

[YKW21] Jan Youtie, Seokbeom Kwon, and Seokkyun Woo. The impact of I-Corps on accelerating venture discontinuation in a southeastern US university. Science and Public Policy, 48 (4):474–487, August 2021.

CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/4/474/6231618.

Yang:2023:REI

[YLGZ23] Boxu Yang, Xielin Liu, Yuchen Gao, and Langmei Zhu. A research on the effectiveness of innovation policy for regional innovation under Chinese long-range plan. Science and Public Policy, 50(3): 491–508, June 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/3/491/7059261.

Yaqub:2022:RDR

[YLMYY22] Ohid Yaqub, Javier A. Luna, Duncan Aq Moore, and Alfredo Yegros-Yegros. Responding to a disease with resources from other diseases: Evidence from Zika vaccine research dynamics. Science and Public Policy, 49(6):942–950, December 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/6/942/6657697.

Yu:2021:EEG

[YXX21] Hanzhi Yu, Jianhua Xu, and Lan Xue. Engaging experts in global biotech governance: What influences their judgement on valueladen challenges? Science and Public Policy, 48(4): 553–561, August 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/48/4/553/6291353.

Zhao:2020:CSM

[ZAB^{+}20] Jiaying Zhao, Meghan B.

Azad, Erin M. Bertrand, Cole Burton, Valorie A. Crooks, Jackie Dawson, Adam T. Ford, Angela Kaida, Arjun Krishnaswamy, Chikin Kuok, Catherine L. Mah, Matt McTaggart, Amanda J. Moehring, Dominique Robert, Albrecht Schulte-Hostedde, Heather Sparling, Mary A. De Vera, Stephanie Waterman, and Trushar R. Patel. Canadian science meets Parliament: Building relationships between scientists and policymakers. Science and Public Policy, 47 (2):298, April 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/2/298/5760335.

Zhao:2021:CSM

[ZAB^{+}21] Jiaying Zhao, Meghan B.

Azad, Erin M. Bertrand, Cole Burton, Valorie A. Crooks, Jackie Dawson, Adam T. Ford, Angela Kaida, Arjun Krishnaswamy, Chikin Kuok, Catherine L. Mah, Matt McTaggart, Amanda J. Moehring, Dominique Robert, Albrecht Schulte-Hostedde, Heather Sparling, Mary A. De Vera, Stephanie Waterman, and Trushar R. Patel. Canadian science meets Parliament: Building relationships between scientists and policymakers. Science and Public Policy, 48(4):447– 450, August 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/48/4/447/5710706.

Zhao:2021:LGC

[ZFCL21] Chen Zhao, Feng Feng, Yuee Chen, and Xingteng Li. Local government competition and regional innovation efficiency: From the perspective of China-style fiscal federalism. Science and Public Policy, 48(4): 488–489, August 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/48/4/488/6238579.

Zhang:2022:HPE

[ZG22] Chao Zhang and Jiancheng Guan. How policies emerge and interact with each other? A bibliometric analysis of policies in China. Science and Public Policy, 49(3):441–459, June 2022.

CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/3/441/6513776.

Zou:2022:BSR

[Zou22] Yawen Zou. A bibliometric study on the R&D funding and academic research performance in Shenzhen. Science and Public Policy, 49(3):460–473, June 2022. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/49/3/460/6523028.

Zhang:2023:CPU

[ZQH^{+}23] Lin Zhang, Fan Qi, Ying Huang, Bart Van Looy, Lixin Chen, and Ozcan Saritas. Chinese public university patents during 2006– 20: a comprehensive investigation and comparative study. Science and Public Policy, 50(3):416– 432, June 2023. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/50/3/416/7042012.

Zavale:2021:ASE

[ZS21] Nelson Casimiro Zavale and Christian Schneijderberg. Academics’ societal engagement in ecologies of knowledge: a case study from Mozambique. Science and Public Policy, 48(1):

37–52, February 2021. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/48/1/37/6026551.

Zhou:2020:IPI

[ZW20] William Chongyang Zhou and Ruicheng Wang. Intellectual property, institutional dynamics, and firm innovation. Science and Public Policy, 47(3):299– 312, June 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/47/3/299/5836010.

Zang:2020:HWD

[ZX20] Leizhen Zang and Feng Xiong. How (when) does technological innovation improve government effectiveness? An empirical investigation with crossnational evidence. Science and Public Policy, 47(1): 103–113, February 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

                                        /academic.oup.com/spp/
                                        article/47/1/103/5609055.

Zhang:2020:ASV

[ZXW^{+}20] Gupeng Zhang, Libin Xiong, Xiao Wang, Jianing Dong, and Hongbo Duan. Artificial selection versus natural selection: Which causes the Matthew effect of science funding allocation in China? Science and Public Policy, 47(3): 434–445, June 2020. CO-DEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/3/434/5828332.

Zhou:2020:IPD

[ZYH^{+}20] Ping Zhou, Lichao Yang, Pingbo Hu, Yang Cheng, and Shenghan Zhou. Interprovincial differences of technical innovation efficiency among farmer professional cooperatives in China. Science and Public Policy, 47(3):410–424, June 2020. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic). URL http:/

/academic.oup.com/spp/
article/47/3/410/5828325.